### v1

# Ore extensions of multiplier Hopf coquasigroups

Abstract In this paper, Ore extensions of multiplier Hopf coquasigroups are studied. Necessary and sufficient conditions for the Ore extension of a multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup are given. Then the isomorphism between two Ore extensions is discussed.
Keywords multiplier Hopf coquasigroup, Ore extension.
Mathematics Subject Classification 16T05 $\cdot$ 16T99
1 Introduction
The Ore extension is a well-known method of ring extensions to construct noncommutative rings and algebras. In Hopf algebra theory, Ore extension were also served as a key tool to create noncommutative and non-cocommutative Hopf algebras, such as Panov [9], Beattie [1, 2] and Nenciu [7, 8].
Over time, Hopf algebras has been evolved into broader frameworks. In the 1990s, Van Daele [10] pioneered the theory of multiplier Hopf algebras, breaking the unit constraint of traditional Hopf algebras and offering new tools for quantum groups and noncommutative algebra. Subsequently, its theory was gradually expanded to duality theory, representation theory, and quantum applications. Meanwhile, the theory of Hopf coquasigroups advanced through the groundbreaking work of Klim and Majid [6]. They introduced weak coassociativity to resolve dual construction challenges in non-coassociative structures. Their work also revealed deep connections with algebraic topology.
Subsequently, some algebraists investigated the feasibility of Ore extensions on these structures (Figure 1). Zhao and Lu [14] initiated research on Ore extensions of multiplier Hopf algebras, while Jiao and Meng [3] extended this to Hopf coquasigroups, proposing Hopf coquasigroup-Ore extensions.
[…]
Figure 1: Diagram of the Ore Extension Structure Relationships
Based on these advances, a natural question arises: Could Ore extensions also hold for multiplier Hopf coquasigroups? This is the main motivation.
This paper is organized as follows. In section 2, we introduce some basic notions: Ore extension, multiplier algebras and multiplier Hopf coquasigroups, which will be used in the following sections.
In section 3, we present the concept of Ore extensions in the context of multiplier Hopf coquasigroups. Moreover, we present the main theorem of this paper (Theorem 3.4), which gives a necessary and sufficient condition for Ore extensions of regular multiplier Hopf coquasigroups to be regular multiplier Hopf coquasigroups. This theorem gives a positive answer to the above question.
In section 4, we give the notion of isomorphism for Ore extensions, and show sufficient conditions for two such extensions to be isomorphic.
2 Preliminaries
Throughout this paper, all linear spaces we considered are over a fixed field $k$ (e.g., the complex number field $\mathds{C}$).
2.1 Multiplier Hopf coquasigroups
Let $A$ be an (associative) algebra over $k$, with or without identity, but the product is nondegenerate. This means that, whenever $a\in A$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, we must have that $a=0$.
Recall from [10, 11] that a left multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho(ab)=\rho(a)b$ for all $a,b\in A$. Similarly, a right multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho(ab)=a\rho(b)$ for all $a,b\in A$. A multiplier of $A$ is a pair $(\rho_{1},\rho_{2})$ of a left and a right multiplier such that $\rho_{2}(a)b=a\rho_{1}(b)$ for all $a,b\in A$. For any algebra $A$, there are a left multiplier algebra, a right multiplier algebra, and a multiplier algebra of $A$, which are respectively denoted by $L(A)$, $R(A)$ and $M(A)$. In particular, $M(A)$ is characterized as the largest algebra with identity containing $A$ as an essential two-sided ideal. Additionally, we still have that, whenever $a\in M(A)$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, again $a=0$. Furthermore, we consider the tensor algebra $A\otimes A$. It is still nondegenerate and we have its multiplier algebra $M(A\otimes A)$. There are natural imbeddings
$A\otimes A\subseteq M(A)\otimes M(A)\subseteq M(A\otimes A).$
In general, when $A$ has no identity, these two inclusions are strict. If $A$ already has an identity $1_{A}$, the product is obviously nondegenerate and $M(A)=A$ and $M(A\otimes A)=A\otimes A$.
Let $A$ and $B$ be nondegenerate algebras, if homomorphism $f:A\longrightarrow M(B)$ is nondegenerate (i.e., $f(A)B=B$ and $Bf(A)=B$), then it has a unique extension to a homomorphism $M(A)\longrightarrow M(B)$, we also denote it $f$.
A regular multiplier Hopf coquasigroup introduced in [13] is a nondegenerate associative algebra $A$ equipped with algebra homomorphisms $\Delta:A\longrightarrow M(A\otimes A)$(coproduct), $\varepsilon:A\longrightarrow k$ (counit), and a linear map $S:A\longrightarrow A$ (antipode) such that
  1. (1)
    $T_{1}(a\otimes b)=\Delta(a)(1\otimes b)$ and $T_{2}(a\otimes b)=(a\otimes 1)\Delta(b)$ belong to $A\otimes A$ for any $a,b\in A$.
  2. (2)
    The counit satisfies $(\varepsilon\otimes id)T_{1}(a\otimes b)=ab=(id\otimes\varepsilon)T_{2}(a\otimes b)$.
  3. (3)
    $S$ is antimultiplicative and anticomultiplicative such that for any $a,b\in A$
    $(m\otimes\iota)(S\otimes\Delta)\Delta(a)=1_{M(A)}\otimes a=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(a),$ (2.1)
    $(\iota\otimes m)(\Delta\otimes S)\Delta(a)=a\otimes 1_{M(A)}=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(a).$ (2.2)
Multiplier Hopf coquasigroup $(A,\Delta)$ is called regular, if the antipode $S$ is bijective. $T_{1}$ and $T_{2}$ in the above are usually called Galois maps.
Multiplier Hopf coquasigroups introduced in[13] generalized Hopf coquasigroups to a non-unital case. Unlike Hopf coquasigroups in [5, 6], the underlying algebra of multiplier Hopf coquasigroup $A$ need not have an identity in general, and the comultiplication $\Delta$ does not map $A$ into $A\otimes A$ but rather into its multiplier algebra $M(A\otimes A)$. Multiplier Hopf coquasigroups weaken the coassociativity of multiplier Hopf algebra, and unify the notions of multiplier Hopf algebras and Hopf coquasigroups.
We use the adapted Sweedler notation introduced in [12] for the comultiplication: $\Delta(a)=a_{(1)}\otimes a_{(2)}$ for $a\in A$.
2.2 Ore extention
Ore extension of an algebra
First, let us recall the definition of Ore extension of an algebra from [4]. Let $A$ be an algebra, and $\tau{}$ be an algebra endomorphism of $A$. A linear endomorphism $\delta{}$ of $A$ is called an $\tau$-derivation $\delta$ of $A$ if for all $a,b\in A$,
$\delta(ab)=\delta(a)b+\tau(a)\delta(b)$
The above condition implies $\delta(1_{A})=0$.
The Ore extension $A[y;\tau,\delta]$ of an algebra $A$ is an algebra generated by the algebra $A$ with the relation
$ya=\tau(a)y+\delta(a)$
for all $a\in A$.
Ore extension of Hopf algebra
Let $A$ and $A[y;\tau,\delta]$ be Hopf algebras over $k$. The Hopf algebra $A[y;\tau,\delta]$ is called the Hopf-Ore extension if $\Delta(y)=y\otimes 1+r\otimes y$ for some $r\in A$ and $A$ is a Hopf subalgebra in $A[y;\tau,\delta]$.
Following Theorem 1.3 in [9], the necessary and sufficient conditions for the Hopf algebra $A[y;\tau,\delta]$ to be a Hopf-Ore extension are given as follows.
(1) there is a character $\chi:A\rightarrow k$ such that
$\tau(a)=\chi(a_{(1)})a_{(2)}\qquad\text{for all $a\in A$};$
(2) the following relation holds:
$\chi(a_{(1)})a_{(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)});$
(3) the $\tau$-derivation $\delta$ satisfies the relation
$\Delta\big{(}\delta(a)\big{)}=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)}).$
Ore extension of Hopf coquasigroups
Let $A$ and $A[y;\tau,\delta]$ be Hopf coquasigroups, recall from [3] that the Hopf coquasigroup $A[y;\tau,\delta]$ is a Hopf coquasigroups-Ore extension if $\Delta(y)=y\otimes 1+r\otimes y$ for some $r\in A$ and $A$ is a sub-Hopf coquasigroup of $A[y;\tau,\delta]$.
From Theorem 3.3 in [3], for a Hopf coquasigroup $A$, the Hopf coquasigroup $H=A[y,\tau,\delta]$ becomes an Hopf coquasigroup-Ore extension if and only if the following conditions hold:
(1) There exists a character $\chi:A\rightarrow k$, such that for every $a\in A$, satisfying
$\alpha(a)=\chi(a_{(1)})a_{(2)}.$
(2) The subsequent relations hold:
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=\text{Ad}_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}.$
(3) For the $\alpha$-derivation $\delta$, the relation
$\Delta(\delta(a))=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})$
holds for all $a\in A$.
Ore extension of Multiplier Hopf algebras
Let $(A,\Delta)$ be a regular multiplier Hopf algebra and $A[y;\tau,\delta]$ be an Ore extension of the algebra $A$.
If $\tau$ is a nondegenerate algebra endomorphism, then $\tau$-derivation $\delta$ uniquely extends to the multiplier algebra $M(A)$ (i.e., $\overline{\delta}:M(A)\rightarrow M(A)$). If $\tau$ is surjective, following Proposition 2.3 of Section 2 in [14], the product of $A[y,\tau,\delta]$ is nondegenerate. There are some natural extensions: the comultiplication $\overline{\Delta}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, the antipode $\overline{S}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$, the counit $\overline{\varepsilon}:M(A)[y;\tau,\delta]\rightarrow k$.
If restricting these extensions to the subspace $A[y;\tau,\delta]$, i.e., $\overline{\Delta}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, $\overline{S}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$, $\overline{\varepsilon}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow k$, we get the comultiplication $\Delta^{\prime}$, counit $\varepsilon^{\prime}$, and the antipode $S^{\prime}$ on the Ore extension $(A[y;\tau,\delta],\Delta^{\prime})$ of $A$.
$(A[y;\tau,\delta],\Delta^{\prime})$ is called a Hopf-Ore extension of the multiplier Hopf algebra $A$, if $(A[y;\tau,\delta],\Delta^{\prime})$ is a regular multiplier Hopf algebra such that $\overline{\Delta}(y)=y\otimes 1+r\otimes y$ for some $r\in M(A)$.
For a regular multiplier Hopf algebra $A$ with Ore extension $A[y;\tau,\delta]$, where $\tau$ is surjective. $A[y;\tau,\delta]$ constitutes a Hopf-Ore extension if and only if there exists a group-like element $r\in M(A)$ satisfying the following equivalent conditions:
(1) There exists a character $\chi:A\rightarrow k$ such that, for any $a\in A$, the following equality holds:
$\tau(a)=(\chi\otimes id)\Delta(a)=r\big{(}(id\otimes\chi)\Delta(a)\big{)}r^{-1}.$
(2) For the $\tau$-derivation $\delta$, for all $a\in A$, the relation
$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes id)\Delta(a)+(r\otimes 1)\big{(}(id\otimes\delta)\Delta(a)\big{)}$
is satisfied.
As introduced in Section 2.1 of this paper, in the following, we also denote these unique extensions $\overline{\Delta}$, $\overline{S}$, $\overline{\varepsilon}$ as $\Delta$, $S$, $\varepsilon$ without confusion.
3 Ore extensions of multiplier Hopf coquasigroups
Let $(A,\Delta)$ be a regular multiplier Hopf coquasigroup introduced in [13]. The aim of this section is to show a criterion for an Ore extension of a regular multiplier Hopf coquasigroup to be a regular multiplier Hopf coquasigroup.
First, let us introduce the definition of Ore extension of regular multiplier Hopf coquasigroups, which genelized the Ore extension of Hopf coquasigroups and multiplier Hopf algebras introduced in [3, 14].
Definition 3.1. Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups. The multiplier Hopf coquasigroup $(A[y;\tau,\delta],\Delta)$ is called the Ore extension of multiplier Hopf coquasigroups (or MHC-Ore extension in short) if $\Delta(y)=y\otimes r_{1}+r_{2}\otimes y$ for some $r_{1},r_{2}\in M(A)$.
Remark 3.2. (1) As in [14], $y\otimes 1+r\otimes y$ in above definition belongs to $M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$.
(2) If the coproduct $\Delta$ of multiplier Hopf coquasigroup $A$ is coassociative, then $A$ is a multiplier Hopf algebra and the Ore extension is actually the one introduced in [14].
(3) If the underlying algebra of multiplier Hopf coquasigroup $A$ has an identity, then $A$ is a Hopf coquasigroup and the Ore extension is actually the one introduced in [3].
By the definition of regular multiplier Hopf coquasigroup, for $y\in A[y;\tau,\delta]$, we have
$(m\otimes\iota)(S\otimes\Delta)\Delta(y)=1_{M(A)}\otimes y=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(y),$
$(\iota\otimes m)(\Delta\otimes S)\Delta(y)=y\otimes 1_{M(A)}=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(y),$
that is
$S(y)r_{1(1)}\otimes r_{1(2)}+S(r_{2})y\otimes r_{1}+S(r_{2})r_{2}\otimes y=1_{M(A)}\otimes y$
$=yS(r_{1(1)})\otimes r_{1(2)}+r_{2}S(y)\otimes r_{1}+r_{2}S(r_{2})\otimes y,$
$y\otimes S(r_{1})r_{1}+r_{2}\otimes S(y)r_{1}+r_{2(1)}\otimes S(r_{2(2)})y=y\otimes 1_{M(A)}$
$=y\otimes r_{1}S(r_{1})+r_{2}\otimes yS(r_{1})+r_{2(1)}\otimes r_{2(2)}S(y).$
Then
$\hskip 40.00006pt{r_{1}}S({r_{1}})=S({r_{1}}){r_{1}}=1,\qquad{r_{2}}S({r_{2}})=S({r_{2}}){r_{2}}=1,$
$S(y){r_{1(1)}}\otimes{r_{1(2)}}+S({r_{2}})y\otimes{r_{1}}=yS({r_{1(1)}})\otimes{r_{1(2)}}+{r_{2}}S(y)\otimes{r_{1}}=0,$
${r_{2}}\otimes S(y){r_{1}}+{r_{2(1)}}\otimes S({r_{2(2)}})y={r_{2}}\otimes yS({r_{1}})+{r_{2(1)}}\otimes{r_{2(2)}}S(y)=0.$
Let $r_{1}$ and $r_{2}$ be group-like elements in $M(A)$, then we have
$\Delta(r_{i})=r_{i}\otimes r_{i},\quad\varepsilon(r_{i})=1,\quad S(r_{i})=r_{i}^{-1},\quad i=1,2,$
$\big{(}S(y)r_{1}+S(r_{2})y\big{)}\otimes r_{1}=\big{(}yS(r_{1})+r_{2}S(y)\big{)}\otimes r_{1}=0,$
$r_{2}\otimes\big{(}S(y)r_{1}+S(r_{2})y\big{)}=r_{2}\otimes\big{(}yS(r_{1})+r_{2}S(y)\big{)}=0.$
By substituting the generator $y$ with $y^{\prime}=yr_{1}^{-1}$, it follows that the coproduct map satisfies $\Delta(y^{\prime})=y^{\prime}\otimes 1+r\otimes y^{\prime}$, where $r=r_{2}r_{1}^{-1}$ is a group-like element in the multiplier algebra $M(A)$. In the subsequent discussion, we consistently assume that the element $y$ in the Ore extension of a regular multiplier Hopf coquasigroup, the following relation holds:
$\Delta\left(y\right)=y\otimes 1+r\otimes y$ (3.1)
for some group-like element $r\in M(A)$. As usual, $A{d_{r}}\left(a\right)=raS\left(r\right)=ra{r^{-1}}$.
Now we give the following properties for the Ore extension of a regular multiplier Hopf coquasigroup.
Proposition 3.3. When $A[y;\tau,\delta]$ is a MHC-Ore extension of $A$, then
$\varepsilon(y)=0,\quad S(y)=-r^{-1}y,\quad S^{-1}(y)=-\tau(r)^{-1}y,$
$\hskip 60.00009pt\delta(r)=0,\quad\delta(r^{-1})=0.$
proof. By applying the main property of counit, where $a=a_{(1)}\varepsilon(a_{(2)})$, it follows that $y=\varepsilon(y)+\varepsilon(r)y=\varepsilon(y)+y$, this implies $\varepsilon(y)=0$.
By applying the main property of antipode, where $\varepsilon(a)=a_{(1)}S(a_{(2)})$, it follows that $0=\varepsilon(y)=y+rS(y)$, this implies $S(y)=-r^{-1}y$. Since
$y=S^{-1}S(y)=S^{-1}(-r^{-1}y)=-S^{-1}(r^{-1}y)=-S^{-1}(y)S^{-1}(r^{-1})=-S^{-1}(y)r,$
there is an element $a\in A$ such that $S^{-1}(y)=ay$. And applying the preceding formula we get
$y=-ayr=-a\tau(r)y-a\delta(r),$
this implies that $a=-\tau(r)^{-1}$ and $\delta(r)=0$, so we get that $S^{-1}(y)=-\tau(r)^{-1}y$. Since $\delta$ is a $\tau-derivation$, we have
$0=\delta(1)=\delta(rr^{-1})=\delta(r)r^{-1}+\tau(r)\delta(r^{-1}),$
[…]
Next, we will present the main result of this section, which gives the necessary and sufficient conditions for the Ore extension of a multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup.
Theoerm 3.4. Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups. Assume that $\tau$ is a surjective endomorphism on $A$. Then the regular multiplier Hopf coquasigroup $A[y;\tau,\delta]$ is a MHC-Ore extension of $A$ if and only if there is a group-like element $r\in M(A)$ such that the following conditions hold:
(1) There exists a character $\chi$: $A\rightarrow k$ satisfying
$\tau(a)=\chi(a_{(1)})a_{(2)};$ (3.2)
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)};$ (3.3)
for any $a\in A$.
(2) The $\tau-derivation$ $\delta$ satisfies the relation
$\Delta(\delta(a))=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})$ (3.4)
for any $a\in A$.
Proof. The proof will be conducted in three steps.
step1. Comultiplication.
Similar to the contruction in [14], assume that the comultiplication $\Delta$ can be extended from $A$ to $A[y;\tau.\delta]$ by $\Delta(y)=y\otimes 1+r\otimes y$. Then the algebra homomorphism $\Delta$ preserves the relation $ya=\tau(a)y+\delta(a)$ for any $a\in A$, i.e.,
$\Delta(y)\Delta(a)=\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}.$
For any $a,b\in A$, we have that
$\Delta(y)\Delta(a)(1\otimes b)$
$=(y\otimes 1+r\otimes y)(a_{(1)}\otimes{a_{(2)}}b)$
$=y{a_{(1)}}\otimes{a_{(2)}}b+r{a_{(1)}}\otimes y{a_{(2)}}b$
$=\tau({a_{(1)}})y\otimes{a_{(2)}}b+\delta({a_{(1)}})\otimes{a_{(2)}}b+r{a_{(1)}}\otimes\tau({a_{(2)}}b)y+r{a_{(1)}}\otimes\delta({a_{(2)}}b)$
$=\tau({a_{(1)}})y\otimes{a_{(2)}}b+\delta({a_{(1)}})\otimes{a_{(2)}}b+r{a_{(1)}}\otimes\tau({a_{(2)}})\tau(b)y+r{a_{(1)}}\otimes\delta({a_{(2)}})b$
$\quad+r{a_{(1)}}\otimes\tau({a_{(2)}})\delta(b)$
$=\big{(}\tau({a_{(1)}})\otimes a_{(2)}\big{)}(1\otimes b)(y\otimes 1)+\big{(}r{a_{(1)}}{r^{-1}}\otimes\tau({a_{(2)}})\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)$
$\quad+\big{(}r{a_{(1)}}{r^{-1}}\otimes\delta({a_{(2)}})\big{)}(r\otimes b)+\big{(}r{a_{(1)}}{r^{-1}}\otimes\tau({a_{(2)}})\big{)}\big{(}r\otimes\delta(b)\big{)}+\big{(}\delta(a_{(1)})\otimes a_{(2)}\big{)}(1\otimes b),$
and
$\Big{(}\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}\Big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(y\otimes b+r\otimes yb)+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(y\otimes b)+\Delta\big{(}\tau(a)\big{)}(r\otimes yb)+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)y\big{)}+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}$
$\quad+\Delta\big{(}\delta(a)\big{)}(1\otimes b).$
So we have that $\Delta(y)\Delta(a)=\Delta(\tau(a))\Delta(y)+\Delta(\delta(a))$ holds if and only if the following formulas hold:
$\Delta\big{(}\tau(a)\big{)}(1\otimes b)$ $=\tau({a_{(1)}})\otimes{a_{(2)}}b,$
$\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)\big{)}$ $={r{a_{(1)}}\otimes\tau({a_{(2)}}b)},$
$\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$ $=r{a_{(1)}}\otimes\delta({a_{(2)}}b)+\delta({a_{(1)}})\otimes{a_{(2)}}b.$
By the definition of $\delta\otimes\iota$ and $\iota\otimes\delta$ in [14, Lemma 2.5] and the extensions of $\tau\otimes\iota$, the above conditions change to the following:
$\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)},$ (3.5)
$\Delta\big{(}\tau(a)\big{)}=A{d_{r}}(a_{(1)})\otimes\tau(a_{(2)}),$ (3.6)
$\Delta\big{(}\delta(a)\big{)}=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)}).$
[…]
$\chi(a)b=m(\iota\otimes S)\big{(}(b\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
for any $a,b\in A$. Because $\tau$ is nondegenerate, for $b$ there are $a_{i}$ , $b_{i}$ such that $\sum{{a_{i}}}\tau\left({{b_{i}}}\right)=b$. Hence, the above definition is changed to the following:
$\chi(a)b$ $=m(id\otimes S)\Big{(}\sum\big{(}a_{i}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}$
$=m(id\otimes S)\Big{(}\sum(a_{i}\otimes 1)\big{(}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}$
$=m(id\otimes S)\bigg{(}\sum(a_{i}\otimes 1)\Big{(}(\tau\otimes id)(b_{i}\otimes 1)\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}\bigg{)}$
$=m(id\otimes S)\bigg{(}\sum(a_{i}\otimes 1)(\tau\otimes id)\Big{(}(b_{i}\otimes 1)\Delta\big{(}\tau(a)\big{)}\Big{)}\bigg{)}.$
Hence $\chi(a)b\in A$ for any $a,b\in A$. In addition, for any $c\in A$,
$\chi(a)(bc)$ $=m(id\otimes S)\big{(}(bc\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
$=m(id\otimes S)\big{(}(b\otimes 1)(c\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
$=b\big{(}m(id\otimes S)(c\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
$=b\big{(}\chi(a)c\big{)}.$
[…]
$\chi(a)b=m(id\otimes S)\Big{(}(b\otimes 1)(\tau\otimes id)\Delta\big{(}\tau(a)\big{)}\Big{)}=b\varepsilon\big{(}\tau(a)\big{)},$
and since $b$ is arbitrary, this proves that $\chi(a)=\varepsilon\big{(}\tau(a)\big{)}.$ Therefore, $\chi(A)\subseteq k$. We can regard $\chi$ as a mapping $\chi:A\rightarrow k$. Considering that $\tau$ is an algebra endomorphism and
$\chi(ab)=\varepsilon\big{(}\tau(ab)\big{)}=\varepsilon\big{(}\tau(a)\big{)}\varepsilon\big{(}\tau(b)\big{)}=\chi(a)\chi(b),$
$\chi(a+b)=\varepsilon\big{(}\tau(a+b)\big{)}=\varepsilon\big{(}\tau(a)\big{)}+\varepsilon\big{(}\tau(b)\big{)}=\chi(a)+\chi(b),$
[…]
$(id\otimes\Delta)\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)(1)}\otimes a_{(2)(2)}.$
[…]
$S\big{(}\tau(a_{(1)})\big{)}a_{(2)(1)}\otimes a_{(2)(2)}=1\otimes\tau(a)=\tau(a_{(1)})S(a_{(2)(1)})\otimes a_{(2)(2)}.$ (3.7)
By applying the formula $\chi\left(a\right)=\varepsilon\left(\tau\left(a\right)\right)$, $\tau$ can be derived from $\chi$,
$\chi(a_{(1)})a_{(2)}$ $=\varepsilon\big{(}\tau(a_{(1)})\big{)}a_{(2)}=\varepsilon\big{(}\tau(a_{(1)})\big{)}\varepsilon(a_{(2)(1)})a_{(2)(2)}$
$=\varepsilon\big{(}\tau(a_{(1)})\big{)}\varepsilon\big{(}S(a_{(2)(1)})\big{)}a_{(2)(2)}$
$=\varepsilon\big{(}\tau(a_{(1)})\big{)}S(a_{(2)(1)})a_{(2)(2)}$
$=\varepsilon(1)\tau(a)=\tau(a).$
[…]
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)},$
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=A{d_{r}}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}.$
[…]
$\Delta\big{(}\tau(a)\big{)}$ $=\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
$=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}=\tau(a_{(1)})\otimes a_{(2)},$
$\Delta\big{(}\tau(a)\big{)}$ $=\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
$=Ad_{r}(a_{(1)})\otimes\chi(a_{(2)(1)})a_{(2)(2)}=Ad_{r}(a_{(1)})\otimes\tau(a_{(2)}).$
[…]
Therefore, we can obtain that the comultiplication of $A$ (i.e., the comultiplication on $M(A)$ ) can be extended to an algebra homomorphism $\Delta:M(A)[y;\tau,\delta]\rightarrow M(A)[y;\tau,\delta]\otimes M(A)[y;\tau,\delta]$, and $\Delta$ does not necessarily have to be coassociative.
Step2. Counit.
Assume that $A[y;\tau,\delta]$ already has the comultiplication defined as in the above step. In what follows, we will prove the existence of a counit denoted by $\overline{\varepsilon}$ extending $\varepsilon$ from $A$(in fact $M(A)$) to the Ore extension $M(A)[y;\tau,\delta]$. Define $\overline{\varepsilon}(y)=0$. We need to verify the equation
$\overline{\varepsilon}(y)\varepsilon(x)=\varepsilon\big{(}\tau(x)\big{)}\overline{\varepsilon}(y)+\varepsilon\big{(}\delta(x)\big{)}$
for any $x\in M(A)$, i.e., $\varepsilon\big{(}\delta(x)\big{)}=0$. In fact, we only have to prove the above formula is true for any $a\in A$ in the sense that
$0=\varepsilon\big{(}\delta(xa)\big{)}=\varepsilon\big{(}\delta(x)\big{)}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\varepsilon\big{(}\delta(x)\big{)}.$
[…]
$\delta(a)$ $=m(id\otimes\varepsilon)\Delta\big{(}\delta(a)\big{)}$
$=m(id\otimes\varepsilon)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
$=\delta(a_{(1)})\varepsilon(a_{(2)})+ra_{(1)}\otimes\varepsilon\big{(}\delta(a_{(2)})\big{)}$
$=\delta(a)+ra_{(1)}\otimes\varepsilon\big{(}\delta(a_{(2)})\big{)}.$
Therefore, $a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}=0$. Finally,
$\varepsilon\big{(}\delta(a)\big{)}=\varepsilon\Big{(}\delta\big{(}\varepsilon(a_{(1)})a_{(2)}\big{)}\Big{)}=\varepsilon(a_{(1)})\varepsilon\big{(}\delta(a_{(2)})\big{)}=\varepsilon\Big{(}a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}\Big{)}=0$
This proves $\varepsilon\big{(}\delta(a)\big{)}=0$ , for any $a\in A$. Furthermore, since
$0=\varepsilon\big{(}\delta(xa)\big{)}$ $=\varepsilon\big{(}\delta(x)a+\tau(x)\delta(a)\big{)}$
$=\varepsilon\big{(}\delta(a)\big{)}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\varepsilon\big{(}\delta(x)\big{)}$
$=\varepsilon\big{(}\delta(x)\big{)}\Big{(}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\Big{)},$
it follows that for any $x\in M(A)$, $\varepsilon\big{(}\delta(x)\big{)}=0$ holds true. Consequently, $\varepsilon$ can be extended to $M(A)[y;\tau,\delta]$.
For all $\sum a_{i}y^{i},\sum b_{j}y^{j}\in A[y;\tau,\delta]$, it is easy to prove
$(\overline{\varepsilon}\otimes\iota)\big{(}\Delta(\sum a_{i}y^{i})(1\otimes\sum b_{j}y^{j})\big{)}$ $=(\sum a_{i}y^{i})(\sum b_{j}y^{j}),$
$(\iota\otimes\overline{\varepsilon})\big{(}(\sum b_{j}y^{j}\otimes 1)\Delta(\sum a_{i}y^{i})\big{)}$ $=(\sum b_{j}y^{j})(\sum a_{i}y^{i}).$
Step3. Antipode.
[…]
$(m\otimes\iota)(S\otimes\Delta)\Delta(ya)=1\otimes ya=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(ya),$ (3.8)
$(\iota\otimes m)(\Delta\otimes S)\Delta(ya)=ya\otimes 1=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(ya),$ (3.9)
and
$S(a)S(y)=S(y)S\big{(}\alpha(a)\big{)}+S\big{(}\delta(a)\big{)}$ (3.10)
[…]
$1\otimes\delta(a)$ $=(m\otimes\iota)(S\otimes\Delta)\Delta\big{(}\delta(a)\big{)}$
$=(m\otimes\iota)(S\otimes\Delta)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
$=(m\otimes\iota)\Big{(}S\big{(}\delta(a_{(1)})\big{)}\otimes\Delta(a_{(2)})+S(ra_{(1)})\otimes\Delta\big{(}\delta(a_{(2)})\big{)}\Big{)}$
$=(m\otimes\iota)\Big{(}S\big{(}\delta({a_{(1)}})\big{)}\otimes a_{(2)(1)}\otimes a_{(2)(2)}+S(ra_{(1)})\otimes\delta(a_{(2)(1)})\otimes a_{(2)(2)}$
$\quad+S(ra_{(1)})\otimes ra_{(2)(1)}\otimes\delta(a_{(2)(2)})\Big{)}$
$=S\big{(}\delta(a_{(1)})\big{)}a_{(2)(1)}\otimes a_{(2)(2)}+S(ra_{(1)})\delta(a_{(2)(1)})\otimes a_{(2)(2)}$
$\quad+S(ra_{(1)})ra_{(2)(1)}\otimes\delta(a_{(2)(2)}).$ (3.11)
Applying the formula mentioned above, we can obtain
$\big{(}S(b)\otimes 1\big{)}(m\otimes\iota)(S\otimes\Delta)\Delta(ya)$
$=(m\otimes\iota)(S\otimes\Delta)\big{(}\Delta(ya)(b\otimes 1)\big{)}$
$=(m\otimes\iota)(S\otimes\Delta)\Big{(}\Delta\big{(}\tau(a)y+\delta(a)\big{)}(b\otimes 1)\Big{)}$
$=(m\otimes\iota)(S\otimes\Delta)\bigg{(}\Big{(}\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}\Big{)}(b\otimes 1)\bigg{)}$
$=(m\otimes\iota)(S\otimes\Delta)\Big{(}\tau({a_{(1)}})yb\otimes{a_{(2)}}+\tau({a_{(1)}})rb\otimes{a_{(2)}}y+\Delta\big{(}\delta(a)\big{)}(b\otimes 1)\Big{)}$
$=(m\otimes\iota)\Big{(}S\big{(}\tau(a_{(1)})yb\big{)}\otimes{a_{(2)(1)}}\otimes{a_{(2)(2)}}+S\big{(}\tau({a_{(1)}})rb\big{)}\otimes{a_{(2)(1)}}y\otimes{a_{(2)(2)}}$
$\quad+S\big{(}\tau({a_{(1)}})rb\big{)}\otimes a_{(2)(1)}r\otimes{a_{(2)(2)}}y\Big{)}$
$\quad+(m\otimes\iota)(S\otimes\Delta)\big{(}\delta({a_{(1)}})b\otimes{a_{(2)}}+r{a_{(1)}}b\otimes\delta({a_{(2)}})\big{)}$
$=S\big{(}\tau({a_{(1)}})yb\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}+S\big{(}\tau({a_{(1)}})rb\big{)}{a_{(2)(1)}}y\otimes{a_{(2)(2)}}$
$\quad+S\big{(}\tau({a_{(1)}})rb\big{)}{a_{(2)(1)}}r\otimes{a_{(2)(2)}}y+(m\otimes\iota)\Big{(}S\big{(}\delta(a_{(1)})b\big{)}\otimes{a_{(2)(1)}}\otimes{a_{(2)(2)}}$
$\quad+S(r{a_{(1)}}b)\otimes\delta({a_{(2)(1)}})\otimes{a_{(2)(2)}}+S(r{a_{(1)}}b)\otimes r{a_{(2)(1)}}\otimes\delta({a_{(2)(2)}})\Big{)}$
$=S(b)S(y)S\big{(}\tau({a_{(1)}})\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}+S(b)S(r)S\big{(}\tau({a_{(1)}})\big{)}{a_{(2)(1)}}y\otimes{a_{(2)(2)}}$
$\quad+S(b)S(r)S\big{(}\tau({a_{(1)}})\big{)}{a_{(2)(1)}}r\otimes{a_{(2)(2)}}y+S\big{(}\delta(a)b\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}$
$\quad+S(r{a_{(1)}}b)\delta({a_{(2)(1)}})\otimes{a_{(2)(2)}}+S(r{a_{(1)}}b)r{a_{(2)(1)}}\otimes\delta({a_{(2)(1)}})$
[…]
$\quad+(S(b)\otimes 1)\Big{(}S\big{(}\delta(a_{(1)})\big{)}{a_{(2)(1)}}\otimes a_{(2)(2)}+S(r{a_{(1)})\delta(a_{(2)(1)}})\otimes{a_{(2)(2)}}$
$\quad+S(ra_{(1)})ra_{(2)(1)}\otimes\delta(a_{(2)(2)})\Big{)}$
=(3.11)${=}\big{(}S(b)S(y)+S(b)S(r)y\big{)}\otimes\tau(a)+S(b)\otimes\tau(a)y+(S(b)\otimes 1)(1\otimes\delta(a))$ )
$=S(b)\otimes\tau(a)y+S(b)\otimes\delta(a)$
$=S(b)\otimes\big{(}\tau(a)y+\delta(a)\big{)}$
$=S(b)\otimes ya$
$=\big{(}S(b)\otimes 1\big{)}(1\otimes ya).$
[…]
$-S(a)r^{-1}y$ $=-r^{-1}yS\big{(}\tau(a)\big{)}+S\big{(}\delta(a)\big{)}$
$=-r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}y-r^{-1}\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}+S\big{(}\delta(a)\big{)},$
for any $a\in M(A).$
[…]
$S(a)r^{-1}=r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)},$ (3.12)
$\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}=rS\big{(}\delta(a)\big{)}.$ (3.13)
for any $a\in M(A).$
[…]
In fact,
$\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}$ $=\tau\Big{(}S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}\Big{)}=\tau\big{(}\chi({a_{(1)}})S({a_{(2)}})\big{)}=\chi({a_{(1)}})\tau\big{(}S({a_{(2)}})\big{)}$
$=\chi({a_{(1)}})A{d_{r}}\Big{(}S(a_{(2)(2)})\chi\big{(}S(a_{(2)(1)})\big{)}\Big{)}$
$=\chi(a_{(1)})S(a_{(2)(1)})Ad_{r}\big{(}S(a_{(2)(2)})\big{)}$
$=Ad_{r}\big{(}S(a)\big{)}$
[…]
$S\big{(}\tau(a)\big{)}=S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}=\chi(a_{(1)})S(a_{(2)}).$
[…]
$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=rS\big{(}\delta(a)\big{)}.$ (3.14)
[…]
Applying $m(\iota\otimes S)$ to the equation
$(b\otimes 1)\Delta\big{(}\delta(a)\big{)}=b\delta(a_{(1)})\otimes a_{(2)}+bra_{(1)}\otimes\delta(a_{(2)}),$
it follows that
$m(\iota\otimes S)\Big{(}(b\otimes 1)\Delta\big{(}\delta(a)\big{)}\Big{)}=b\delta(a_{(1)})S(a_{(2)})+bra_{(1)}S(\delta\big{(}{a_{(2)}})\big{)}.$
By applying the properties of the antipode, we have
$m(\iota\otimes S)\Big{(}(b\otimes 1)\Delta\big{(}\delta(a)\big{)}\Big{)}=b\varepsilon\big{(}\delta(a)\big{)}=0,$
i.e,
$bra_{(1)}S\big{(}\delta(a_{(2)})\big{)}=-b\delta(a_{(1)})S(a_{(2)})\qquad\text{for any $a,b\in A$}.$
And considering the nondegeneracy of the product of $A$, this means that the following equation holds for any $a\in A$
$a_{(1)}S\big{(}\delta(a_{(2)})\big{)}=-r^{-1}\delta(a_{(1)})S(a_{(2)}).$ (3.15)
Hence, we have
$rS\big{(}\delta(a)\big{)}$ $=rS(a_{(1)})a_{(2)(1)}S\big{(}\delta(a_{(2)(2)})\big{)}$
=(3.15)${=}-rS(a_{(1)})r^{-1}\delta(a_{(2)(1)})S(a_{(2)(2)})$ )
$=-Ad_{r}S({{a_{(1)}}})\delta(a_{(2)(1)})S(a_{(2)(2)}),$ (3.16)
for any $a\in A$.
In addition, from the action of $\delta$ on the element $\varepsilon(a)=a_{1}S(a_{(2)})$, it follows that
$0=\delta(a_{(1)})S(a_{(2)})+\tau(a_{(1)})\delta\big{(}S(a_{(2)})\big{)},$
i.e.,
$\tau(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=-\delta(a_{(1)})S(a_{(2)}).$ (3.17)
[…]
$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}$ $=\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})a_{(2)(2)(1)}\big{)}\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
$=\chi(a_{(1)})\tau\big{(}{S(a_{(2)(1)}})\big{)}\tau(a_{(2)(2)(1)})\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
=(3.17)${=}-\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})\big{)}\delta(a_{(2)(2)(1)})S(a_{(2)(2)(2)})$ )
$=-\tau\Big{(}S\big{(}\chi(a_{(1)})a_{(2)(1)}\big{)}\Big{)}\delta(a_{(2)(2)(1)})S(a_{(2)(2)(2)})$
=(3.3)${=}-\tau\Big{(}S\big{(}\chi(a_{(1)(1)})a_{(1)(2)}\big{)}\Big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$ )
$=-\chi(a_{(1)(1)})\tau\big{(}S(a_{(1)(2)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
$=-\chi(a_{(1)(1)})Ad_{r}\big{(}S(a_{(1)(2)(2)})\big{)}\chi\big{(}S(a_{(1)(2)(1)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
$=-\chi(a_{(1)(1)})S(a_{(1)(2)(1)})Ad_{r}\big{(}S(a_{(1)(2)(2)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
$=-Ad_{r}\big{(}S(a_{(1)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
=(3.16)${=}rS\big{(}\delta(a)\big{)}$ )
for any $a,b\in A$.
[…]
4 Isomorphism of Ore extensions
From Theorem 3.4, it follows that a MHC-Ore extension is determined by $\chi$, $r$, and $\delta$. Thus, $A(\chi,r,\delta)$ is used to denote the MHC-Ore extension of $A[y;\tau,\delta]$, where $\chi$ is a character of $A$, $r$ is a group-like element in $M(A)$ and $\delta$ is a $\langle\chi,r\rangle$-derivation. The corresponding Ore extension on the multiplier coquasigroup is denoted by $M(A)(\chi,r,\delta)$.
Now we can define an isomorphism of MHC-Ore extensions.
Definition 4.1. Two MHC-Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ of regular multiplier Hopf coquasigroups $A$ and $A^{\prime}$ are said to be isomorphic if there is an isomorphism of multiplier Hopf coquasigroups $\phi$: $A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ such that $\phi(A)=A^{\prime}$.
Similarly to [14, Theorem 4.9], we have the following theorem for regular multiplier Hopf coquasigroups.
Theorem 4.2. For multiplier Hopf coquasigroups $A$ and $A^{\prime}$, let their Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ be defined with coproducts: $\Delta(y)=y\otimes 1+r\otimes y$, $\Delta(y^{\prime})=y^{\prime}\otimes 1+r^{\prime}\otimes y^{\prime}$, where $r$ and $r^{\prime}$ are group-like elements in $A$. Then $A(\chi,r,\delta)$ is isomorphic to $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ if there is an isomorphism $\phi:A\rightarrow A^{\prime}$ such that
$\phi(r)=r^{\prime},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},$ (4.1)
where $d^{\prime}\in A^{\prime}$ such that $\Delta(d^{\prime})=d^{\prime}\otimes 1+r^{\prime}\otimes d^{\prime}$.
Proof. Let $\phi(y)=y^{\prime}+d^{\prime}$. Then $\phi$ can be extended from $\phi:A\rightarrow A^{\prime}$ to $\phi:A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$. We have
$\phi(ya)$ $=\phi\big{(}\tau(a)y+\delta(a)\big{)}=\phi\big{(}\tau(a)\big{)}\phi\big{(}y\big{)}+\phi\big{(}\delta(a)\big{)}$
$=\phi\big{(}\tau(a)\big{)}(y^{\prime}+d^{\prime})+\phi\big{(}\delta(a)\big{)},$
$\phi(y)\phi(a)$ $=(y^{\prime}+d^{\prime})\phi(a)=\tau^{\prime}\big{(}\phi(a)\big{)}y^{\prime}+\delta^{\prime}\big{(}\phi(a)\big{)}+d^{\prime}\phi(a)$
$=\tau^{\prime}\big{(}\phi(a)\big{)}y^{\prime}+\phi\big{(}\delta(a)\big{)}+\phi\big{(}\tau(a)\big{)}d^{\prime}-d^{\prime}\phi(a)+d^{\prime}\phi(a)$
$=\phi\big{(}\tau(a)\big{)}y^{\prime}+\phi\big{(}\tau(a)\big{)}d^{\prime}+\phi\big{(}\delta(a)\big{)}=\phi\big{(}\tau(a)\big{)}(y^{\prime}+d^{\prime})+\phi\big{(}\delta(a)\big{)},$
for any $a\in A$. Thus we get
$\phi(ya)=\phi(y)\phi(a).$
Since
$\Delta\big{(}\phi(y)\big{)}$ $=\Delta(y^{\prime}+d^{\prime})=\Delta(y^{\prime})+\Delta(d^{\prime})$
$=y^{\prime}\otimes 1+r^{\prime}\otimes y^{\prime}+d^{\prime}\otimes 1+r^{\prime}\otimes d^{\prime}$
$=(y^{\prime}+d^{\prime})\otimes 1+r^{\prime}\otimes(y^{\prime}+d^{\prime}),$
$(\phi\otimes\phi)\Delta(y)$ $=(\phi\otimes\phi)(y\otimes 1+r\otimes y)$
$=(y^{\prime}+d^{\prime})\otimes 1+r^{\prime}\otimes(y^{\prime}+d^{\prime}),$
we can obtain
$\Delta\big{(}\phi(y)\big{)}=(\phi\otimes\phi)\Delta(y).$
Considering the properties of $\varepsilon$, we obtain
$d^{\prime}=(\varepsilon\otimes id)\big{(}\Delta(d^{\prime})\big{)}=\varepsilon(d^{\prime})+\varepsilon(r^{\prime})d^{\prime}=\varepsilon(d^{\prime})+d^{\prime},$
this indicates $\varepsilon(d^{\prime})=0$. And by Proposition 3.3, we know that $\varepsilon(y)=0$ and $\varepsilon(y^{\prime})=0$. So we have
$\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y^{\prime}+d^{\prime})=\varepsilon(y^{\prime})+\varepsilon(d^{\prime})=0,$
this means that $\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y)$.
Considering the properties of $S$, we obtain
$0=\varepsilon(d^{\prime})=m(S\otimes id)\Delta(d^{\prime})=S(d^{\prime})+S(r^{\prime})d^{\prime}=S(d^{\prime})+r{{}^{\prime-1}}d^{\prime},$
this indicates $S(d^{\prime})=-r{{}^{\prime-1}}d^{\prime}$. Thus we have
$\phi\big{(}S(y)\big{)}=\phi(-{r^{-1}}y)=-\phi({r^{-1}})\phi(y)=-r{{}^{\prime-1}}(y^{\prime}+d^{\prime}),$
$S\big{(}\phi(y)\big{)}=S(y^{\prime}+d^{\prime})=S(y^{\prime})+S(d^{\prime})=-{r^{-1}}y^{\prime}-r{{}^{\prime-1}}d^{\prime}=-r{{}^{\prime-1}}(y^{\prime}+d^{\prime}),$
it follows that $\phi\big{(}S(y)\big{)}=S\big{(}\phi(y)\big{)}$.
[…]
Acknowledgements
The work was partially supported by the China Postdoctoral Science Foundation (No. 2019M651764) and National Natural Science Foundation of China (No. 11601231).

### v2

@@
 Abstract In this paper, Ore extensions of multiplier Hopf coquasigroups are studied.
-Necessary and sufficient conditions for the Ore extension of a multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup are given.
-Then the isomorphism between two Ore extensions is discussed.
+Necessary and sufficient conditions for the Ore extension of a regular multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup are given.
+Furthermore, the isomorphism between two such Ore extensions is discussed.
 Keywords multiplier Hopf coquasigroup, Ore extension.
@@
 1 Introduction
-The Ore extension is a well-known method of ring extensions to construct noncommutative rings and algebras.
-In Hopf algebra theory, Ore extension were also served as a key tool to create noncommutative and non-cocommutative Hopf algebras, such as Panov [9], Beattie [1, 2] and Nenciu [7, 8].
-Over time, Hopf algebras has been evolved into broader frameworks.
-In the 1990s, Van Daele [10] pioneered the theory of multiplier Hopf algebras, breaking the unit constraint of traditional Hopf algebras and offering new tools for quantum groups and noncommutative algebra.
-Subsequently, its theory was gradually expanded to duality theory, representation theory, and quantum applications.
-Meanwhile, the theory of Hopf coquasigroups advanced through the groundbreaking work of Klim and Majid [6].
-They introduced weak coassociativity to resolve dual construction challenges in non-coassociative structures.
-Their work also revealed deep connections with algebraic topology.
-Subsequently, some algebraists investigated the feasibility of Ore extensions on these structures (Figure 1).
-Zhao and Lu [14] initiated research on Ore extensions of multiplier Hopf algebras, while Jiao and Meng [3] extended this to Hopf coquasigroups, proposing Hopf coquasigroup-Ore extensions.
-[…]
+Hopf algebras have evolved into increasingly general formalisms over the past decades.
+In the early 1990s, Van Daele [10] introduced the concept of multiplier Hopf algebras, a framework that dispenses with the unitality requirement intrinsic to classical Hopf algebras and thereby furnishes new tools for the study of quantum groups and non-commutative algebras.
+This theory has since been systematically expanded to include an intrinsic duality theory, a comprehensive representation theory, and applications in quantum field theory.
+Hopf coquasigroups have undergone substantial development, particularly through the foundational work in 2010 of Klim and Majid [6], who formulated a weak form of coassociativity.
+Their approach establishes profound links with algebraic topology and opens new avenues for the study of non-(co)associative symmetries.
+The Ore extension is a well-known method in ring extensions used to construct noncommutative rings and algebras.
+Moreover, within the theory of Hopf algebras, Ore extensions are significant for constructing noncommutative and non-cocommutative Hopf algebras, as demonstrated in the works of Panov [9], Beattie [1, 2], and Nenciu [7, 8].
+Subsequently, researchers investigated the feasibility of endowing these generalized Hopf-type structures with Ore extensions (see Figure 1).
+Zhao and Lu [16] initiated the systematic study of Ore extensions in the context of multiplier Hopf algebras, while Jiao and Meng [3] formulated Ore extensions for Hopf coquasigroups and established the notion of Hopf coquasigroup–Ore extensions.
+Ore Extensions ofHopf algebrasOre Extensions ofHopf coquasigroupsOre Extensions ofmultiplier Hopf algebrasOre Extensions ofmultiplier Hopf coquasigroupsJiao & Meng[9]Zhao & Lu[16] ??
 Figure 1: Diagram of the Ore Extension Structure Relationships
-Based on these advances, a natural question arises: Could Ore extensions also hold for multiplier Hopf coquasigroups? This is the main motivation.
-This paper is organized as follows. In section 2, we introduce some basic notions: Ore extension, multiplier algebras and multiplier Hopf coquasigroups, which will be used in the following sections.
-In section 3, we present the concept of Ore extensions in the context of multiplier Hopf coquasigroups.
-Moreover, we present the main theorem of this paper (Theorem 3.4), which gives a necessary and sufficient condition for Ore extensions of regular multiplier Hopf coquasigroups to be regular multiplier Hopf coquasigroups.
-This theorem gives a positive answer to the above question.
-In section 4, we give the notion of isomorphism for Ore extensions, and show sufficient conditions for two such extensions to be isomorphic.
+In 2022, Yang [14] established a unifying framework for multiplier Hopf algebras and Hopf coquasigroups through the introduction of multiplier Hopf coquasigroups.
+This theoretical advancement not only bridged these two important algebraic structures but also provided a solution to the biduality problem for a class of infinite-dimensional Hopf quasigroups.
+Motivated by these recent advances, the present work addresses the following fundamental question: Can Ore extensions be realized within the framework of multiplier Hopf coquasigroups?
+This paper provides a positive answer, and its structure is as follows.
+In Section 2, we provide an introduction to key concepts: Ore extensions, multiplier algebras, and multiplier Hopf coquasigroups, which will be referenced in subsequent sections.
+Section 3 discusses Ore extensions within the framework of multiplier Hopf coquasigroups.
+Additionally, one of the main results of this study (Theorem 3.4) is presented, offering a definitive condition for Ore extensions of regular multiplier Hopf coquasigroups to also be regular.
+This theorem affirms the previous question.
+Meanwhile, we consider the *-algebra version.
+Moving on to Section 4, we discuss the isomorphism for Ore extensions and establish the criteria under which two such extensions are considered isomorphic.
 2 Preliminaries
-Throughout this paper, all linear spaces we considered are over a fixed field $k$ (e.g., the complex number field $\mathds{C}$).
+Throughout this paper, all linear spaces under consideration are defined over a fixed field $k$ (for example, the field of complex numbers $\mathds{C}$).
 2.1 Multiplier Hopf coquasigroups
-Let $A$ be an (associative) algebra over $k$, with or without identity, but the product is nondegenerate.
-This means that, whenever $a\in A$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, we must have that $a=0$.
-Recall from [10, 11] that a left multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho(ab)=\rho(a)b$ for all $a,b\in A$.
-Similarly, a right multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho(ab)=a\rho(b)$ for all $a,b\in A$.
-A multiplier of $A$ is a pair $(\rho_{1},\rho_{2})$ of a left and a right multiplier such that $\rho_{2}(a)b=a\rho_{1}(b)$ for all $a,b\in A$.
-For any algebra $A$, there are a left multiplier algebra, a right multiplier algebra, and a multiplier algebra of $A$, which are respectively denoted by $L(A)$, $R(A)$ and $M(A)$.
-In particular, $M(A)$ is characterized as the largest algebra with identity containing $A$ as an essential two-sided ideal.
-Additionally, we still have that, whenever $a\in M(A)$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, again $a=0$.
-Furthermore, we consider the tensor algebra $A\otimes A$.
-It is still nondegenerate and we have its multiplier algebra $M(A\otimes A)$.
-There are natural imbeddings
+Let $A$ be an (associative) algebra over $k$, with or without an identity element, but with a nondegenerate product.
+This condition implies that if $a\in A$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, then $a$ must be equal to 0.
+Recall from [10, 11, 13] that a left multiplier of $A$ is a linear map $\rho_{1}:A\rightarrow A$ satisfying $\rho_{1}(ab)=\rho_{1}(a)b$ for all $a,b\in A$.
+Similarly, a right multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho_{2}(ab)=a\rho_{2}(b)$ for all $a,b\in A$.
+A multiplier of AA consists of a pair $(\rho_{1},\rho_{2})$ comprising a left and a right multiplier, where $\rho_{2}(a)b=a\rho_{1}(b)$ holds for all $a,b\in A$.
+For any algebra AA, there exist a left multiplier algebra, a right multiplier algebra, and a multiplier algebra of $A$, denoted as $L(A)$, $R(A)$, and $M(A)$, respectively.
+Specifically, $M(A)$ is characterized as the largest algebra with identity containing AA as an essential two-sided ideal.
+Moreover, in the case where $x\in M(A)$ and $xa=0$ for all $a\in A$ or $ax=0$ for all $a\in A$, it follows that $x=0$.
+Additionally, we examine the tensor algebra $A\otimes A$, which remains nondegenerate, and we have its multiplier algebra $M(A\otimes A)$.
+There are natural embeddings
 $A\otimes A\subseteq M(A)\otimes M(A)\subseteq M(A\otimes A).$
-In general, when $A$ has no identity, these two inclusions are strict. If $A$ already has an identity $1_{A}$, the product is obviously nondegenerate and $M(A)=A$ and $M(A\otimes A)=A\otimes A$.
-Let $A$ and $B$ be nondegenerate algebras, if homomorphism $f:A\longrightarrow M(B)$ is nondegenerate (i.e., $f(A)B=B$ and $Bf(A)=B$), then it has a unique extension to a homomorphism $M(A)\longrightarrow M(B)$, we also denote it $f$.
-A regular multiplier Hopf coquasigroup introduced in [13] is a nondegenerate associative algebra $A$ equipped with algebra homomorphisms $\Delta:A\longrightarrow M(A\otimes A)$(coproduct), $\varepsilon:A\longrightarrow k$ (counit), and a linear map $S:A\longrightarrow A$ (antipode) such that
+Generally, when AA lacks an identity element, these two embeddings are strict.
+In the scenario where AA possesses an identity element $1_{A}$, the product is evidently nondegenerate, and we observe that $M(A)=A$ and $M(A\otimes A)=A\otimes A$.
+Let $A$ and $B$ be nondegenerate algebras.
+If a homomorphism $f:A\longrightarrow M(B)$ is nondegenerate (meaning $f(A)B=B$ and $Bf(A)=B$), then it can be uniquely extended to a homomorphism $M(A)\longrightarrow M(B)$, also denoted by $f$.
+A multiplier Hopf coquasigroup, as defined in [14], is a nondegenerate associative algebra AA equipped with algebra homomorphisms $\Delta:A\longrightarrow M(A\otimes A)$ (comultiplication), $\varepsilon:A\longrightarrow k$ (counit), and a linear map $S:A\longrightarrow A$ (antipode).
+The key conditions for this structure are as follows:
   1. (1)
-    $T_{1}(a\otimes b)=\Delta(a)(1\otimes b)$ and $T_{2}(a\otimes b)=(a\otimes 1)\Delta(b)$ belong to $A\otimes A$ for any $a,b\in A$.
+    For any $a,b\in A$, both $T_{1}(a\otimes b)=\Delta(a)(1\otimes b)$ and $T_{2}(a\otimes b)=(a\otimes 1)\Delta(b)$ belong to $A\otimes A$.
   2. (2)
-    The counit satisfies $(\varepsilon\otimes id)T_{1}(a\otimes b)=ab=(id\otimes\varepsilon)T_{2}(a\otimes b)$.
+    The counit satisfies $(\varepsilon\otimes\iota)T_{1}(a\otimes b)=ab=(\iota\otimes\varepsilon)T_{2}(a\otimes b)$, where $\iota$ denotes the identity map, and we will use this notation henceforth.
   3. (3)
-    $S$ is antimultiplicative and anticomultiplicative such that for any $a,b\in A$
+    The antipode $S$ is defined to be antimultiplicative and anticomultiplicative, such that for any $a,b\in A$ the following hold:
     $(m\otimes\iota)(S\otimes\Delta)\Delta(a)=1_{M(A)}\otimes a=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(a),$ (2.1)
     $(\iota\otimes m)(\Delta\otimes S)\Delta(a)=a\otimes 1_{M(A)}=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(a).$ (2.2)
-Multiplier Hopf coquasigroup $(A,\Delta)$ is called regular, if the antipode $S$ is bijective. $T_{1}$ and $T_{2}$ in the above are usually called Galois maps.
-Multiplier Hopf coquasigroups introduced in[13] generalized Hopf coquasigroups to a non-unital case.
-Unlike Hopf coquasigroups in [5, 6], the underlying algebra of multiplier Hopf coquasigroup $A$ need not have an identity in general, and the comultiplication $\Delta$ does not map $A$ into $A\otimes A$ but rather into its multiplier algebra $M(A\otimes A)$.
-Multiplier Hopf coquasigroups weaken the coassociativity of multiplier Hopf algebra, and unify the notions of multiplier Hopf algebras and Hopf coquasigroups.
-We use the adapted Sweedler notation introduced in [12] for the comultiplication: $\Delta(a)=a_{(1)}\otimes a_{(2)}$ for $a\in A$.
-2.2 Ore extention
+In the context of a multiplier Hopf coquasigroup $(A,\Delta)$, it is termed regular if the antipode $S$ is bijective. The maps $T_{1}$ and $T_{2}$ are often called the Galois maps.
+Multiplier Hopf coquasigroups, as introduced in [14], extend the concept of Hopf coquasigroups to a non-unital scenario.
+In contrast to the Hopf coquasigroups discussed in [5, 6], the algebraic structure of a multiplier Hopf coquasigroup AA may not necessarily possess an identity element in general.
+Moreover, the comultiplication map $\Delta$ does not merely map elements from $A$ to $A\otimes A$, but rather it maps them into the multiplier algebra $M(A\otimes A)$.
+This extension weakens the coassociativity property typically found in multiplier Hopf algebras while also harmonizing the concepts of multiplier Hopf algebras and Hopf coquasigroups.
+To denote the comultiplication, we employ the adapted Sweedler notation as introduced in [12]: $\Delta(a)=a_{(1)}\otimes a_{(2)}$ for $a\in A$.
+2.2 Ore extension
 Ore extension of an algebra
-First, let us recall the definition of Ore extension of an algebra from [4].
-Let $A$ be an algebra, and $\tau{}$ be an algebra endomorphism of $A$.
-A linear endomorphism $\delta{}$ of $A$ is called an $\tau$-derivation $\delta$ of $A$ if for all $a,b\in A$,
-$\delta(ab)=\delta(a)b+\tau(a)\delta(b)$
-The above condition implies $\delta(1_{A})=0$.
-The Ore extension $A[y;\tau,\delta]$ of an algebra $A$ is an algebra generated by the algebra $A$ with the relation
+We start by recalling the definition of the Ore extension of an algebra, as introduced in [4].
+Given an algebra AA with an identity element $1_{A}$ and an algebra endomorphism $\tau$ of $A$, a linear endomorphism $\delta$ of $A$ is referred to as a $\tau$-derivation if for all $a,b\in A$,
+$\delta(ab)=\delta(a)b+\tau(a)\delta(b).$
+Moreover, it is a consequence of the above condition that $\delta(1_{A})=0$.
+The Ore extension $A[y;\tau,\delta]$ of an algebra AA is constructed as an algebra that extends AA with the defining relation
 $ya=\tau(a)y+\delta(a)$
 for all $a\in A$.
-Ore extension of Hopf algebra
-Let $A$ and $A[y;\tau,\delta]$ be Hopf algebras over $k$.
-The Hopf algebra $A[y;\tau,\delta]$ is called the Hopf-Ore extension if $\Delta(y)=y\otimes 1+r\otimes y$ for some $r\in A$ and $A$ is a Hopf subalgebra in $A[y;\tau,\delta]$.
-Following Theorem 1.3 in [9], the necessary and sufficient conditions for the Hopf algebra $A[y;\tau,\delta]$ to be a Hopf-Ore extension are given as follows.
-(1) there is a character $\chi:A\rightarrow k$ such that
+Ore extension of Hopf algebras
+Let $A$ and $A[y;\tau,\delta]$ denote Hopf algebras over $k$.
+The Hopf algebra $A[y;\tau,\delta]$ is termed a Hopf-Ore extension if there exists an element $r\in A$ such that $\Delta(y)=y\otimes 1+r\otimes y$, and $A$ is a Hopf subalgebra of $A[y;\tau,\delta]$.
+In accordance with Theorem 1.3 in [9], the conditions necessary and sufficient for the Hopf algebra $A[y;\tau,\delta]$ to qualify as a Hopf-Ore extension are as follows:
+$(1)$ There exists a character $\chi:A\rightarrow k$ satisfying
 $\tau(a)=\chi(a_{(1)})a_{(2)}\qquad\text{for all $a\in A$};$
-(2) the following relation holds:
+$(2)$ The following condition holds
 $\chi(a_{(1)})a_{(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)});$
-(3) the $\tau$-derivation $\delta$ satisfies the relation
+$(3)$ The $\tau$-derivation $\delta$ fulfills the relation
 $\Delta\big{(}\delta(a)\big{)}=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)}).$
+Ore extension of Multiplier Hopf algebras
+By generalizing the Ore extension of Hopf algebras to the non-unital case, Zhao and Lu [16] introduced the Ore extension of multiplier Hopf algebras.
+For more details, let $(A,\Delta)$ be a regular multiplier Hopf algebra, and consider $A[y;\tau,\delta]$, an Ore extension of the algebra AA.
+When $\tau$ is a nondegenerate algebra endomorphism, the $\tau$-derivation $\delta$ can be uniquely extended to the multiplier algebra $M(A)$, denoted as $\overline{\delta}:M(A)\rightarrow M(A)$.
+In this case,
+$A[y;\tau,\delta]\subseteq M(A)[y;\tau,\delta]\subseteq M(A[y;\tau,\delta]).$
+Where $\tau$ is surjective, according to Proposition 2.3 of Section 2 in [16], the product of $A[y;\tau,\delta]$ is nondegenerate.
+Zhao and Lu first constructed $M(A)[y;\tau,\delta]$, an Ore extension of $M(A)$.
+The comultiplication, counit, and antipode of AA can be extended to $M(A)[y;\tau,\delta]$: $\overline{\Delta}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, $\overline{\varepsilon}:M(A)[y;\tau,\delta]\rightarrow k$, $\overline{S}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$.
+By restricting these extensions to the subspace $A[y;\tau,\delta]$, the authors obtained the comultiplication, counit, and antipode on the Ore extension $A[y;\tau,\delta]$ of AA:
+$\overline{\Delta}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta]),$
+$\overline{\varepsilon}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow k,$
+$\overline{S}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]).$
+These maps were respectively denoted by $\Delta^{\prime}$, $\varepsilon^{\prime}$, and $S^{\prime}$.
+The pair $(A[y;\tau,\delta],\Delta^{\prime})$ is called a Hopf-Ore extension of the multiplier Hopf algebra AA, if $(A[y;\tau,\delta],\Delta^{\prime},\varepsilon^{\prime},S^{\prime})$ is a regular multiplier Hopf algebra such that $\overline{\Delta}(y)=y\otimes 1_{M(A)}+r\otimes y$ for some $r\in M(A)$, where $1_{M(A)}$ is the identity element of $M(A)$.
+For a regular multiplier Hopf algebra AA with an Ore extension $A[y;\tau,\delta]$, where $\tau$ is surjective, $A[y;\tau,\delta]$ forms a Hopf-Ore extension if and only if there exists a group-like element $r\in M(A)$ satisfying the following equivalent conditions:
+$(1)$ There exists a character $\chi:A\rightarrow k$ such that, for any $a\in A$, the following equality holds:
+$\tau(a)=(\chi\otimes\iota)\Delta(a)=r\big{(}(\iota\otimes\chi)\Delta(a)\big{)}r^{-1}.$
+$(2)$ For the $\tau$-derivation $\delta$, for all $a\in A$, the relation
+$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes\iota)\Delta(a)+(r\otimes 1_{M(A)})\big{(}(\iota\otimes\delta)\Delta(a)\big{)}$
+is satisfied.
+As described in Section 2.1 of this paper, we will refer to these unique extensions as $\overline{\Delta}$, $\overline{S}$, $\overline{\varepsilon}$ simply as $\Delta$, $S$, $\varepsilon$ in the following without causing any confusion.
 Ore extension of Hopf coquasigroups
-Let $A$ and $A[y;\tau,\delta]$ be Hopf coquasigroups, recall from [3] that the Hopf coquasigroup $A[y;\tau,\delta]$ is a Hopf coquasigroups-Ore extension if $\Delta(y)=y\otimes 1+r\otimes y$ for some $r\in A$ and $A$ is a sub-Hopf coquasigroup of $A[y;\tau,\delta]$.
-From Theorem 3.3 in [3], for a Hopf coquasigroup $A$, the Hopf coquasigroup $H=A[y,\tau,\delta]$ becomes an Hopf coquasigroup-Ore extension if and only if the following conditions hold:
-(1) There exists a character $\chi:A\rightarrow k$, such that for every $a\in A$, satisfying
-$\alpha(a)=\chi(a_{(1)})a_{(2)}.$
-(2) The subsequent relations hold:
+Let $A$ and $A[y;\tau,\delta]$ denote Hopf coquasigroups.
+As defined in [3], the Hopf coquasigroup $A[y;\tau,\delta]$ is classified as a Hopf coquasigroups-Ore extension when $\Delta(y)=y\otimes 1+r\otimes y$ for a certain $r\in A$, and $A$ is identified as a sub-Hopf coquasigroup of $A[y;\tau,\delta]$.
+According to Theorem 3.3 in [3], a Hopf coquasigroup AA transforms into a Hopf coquasigroup-Ore extension denoted by $H=A[y,\tau,\delta]$ if and only if the following conditions are met:
+$(1)$ A character $\chi:A\rightarrow k$ exists such that for each $a\in A$, the condition $\tau(a)=\chi(a_{(1)})a_{(2)}$ is satisfied.
+(2)(2)The subsequent relations are valid:
 $\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=\text{Ad}_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}.$
-(3) For the $\alpha$-derivation $\delta$, the relation
+$(3)$ For the $\tau$-derivation $\delta$, the relation
 $\Delta(\delta(a))=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})$
 holds for all $a\in A$.
-Ore extension of Multiplier Hopf algebras
-Let $(A,\Delta)$ be a regular multiplier Hopf algebra and $A[y;\tau,\delta]$ be an Ore extension of the algebra $A$.
-If $\tau$ is a nondegenerate algebra endomorphism, then $\tau$-derivation $\delta$ uniquely extends to the multiplier algebra $M(A)$ (i.e., $\overline{\delta}:M(A)\rightarrow M(A)$).
-If $\tau$ is surjective, following Proposition 2.3 of Section 2 in [14], the product of $A[y,\tau,\delta]$ is nondegenerate.
-There are some natural extensions: the comultiplication $\overline{\Delta}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, the antipode $\overline{S}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$, the counit $\overline{\varepsilon}:M(A)[y;\tau,\delta]\rightarrow k$.
-If restricting these extensions to the subspace $A[y;\tau,\delta]$, i.e., $\overline{\Delta}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, $\overline{S}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$, $\overline{\varepsilon}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow k$, we get the comultiplication $\Delta^{\prime}$, counit $\varepsilon^{\prime}$, and the antipode $S^{\prime}$ on the Ore extension $(A[y;\tau,\delta],\Delta^{\prime})$ of $A$.
-$(A[y;\tau,\delta],\Delta^{\prime})$ is called a Hopf-Ore extension of the multiplier Hopf algebra $A$, if $(A[y;\tau,\delta],\Delta^{\prime})$ is a regular multiplier Hopf algebra such that $\overline{\Delta}(y)=y\otimes 1+r\otimes y$ for some $r\in M(A)$.
-For a regular multiplier Hopf algebra $A$ with Ore extension $A[y;\tau,\delta]$, where $\tau$ is surjective.
-$A[y;\tau,\delta]$ constitutes a Hopf-Ore extension if and only if there exists a group-like element $r\in M(A)$ satisfying the following equivalent conditions:
-(1) There exists a character $\chi:A\rightarrow k$ such that, for any $a\in A$, the following equality holds:
-$\tau(a)=(\chi\otimes id)\Delta(a)=r\big{(}(id\otimes\chi)\Delta(a)\big{)}r^{-1}.$
-(2) For the $\tau$-derivation $\delta$, for all $a\in A$, the relation
-$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes id)\Delta(a)+(r\otimes 1)\big{(}(id\otimes\delta)\Delta(a)\big{)}$
-is satisfied.
-As introduced in Section 2.1 of this paper, in the following, we also denote these unique extensions $\overline{\Delta}$, $\overline{S}$, $\overline{\varepsilon}$ as $\Delta$, $S$, $\varepsilon$ without confusion.
 3 Ore extensions of multiplier Hopf coquasigroups
-Let $(A,\Delta)$ be a regular multiplier Hopf coquasigroup introduced in [13].
-The aim of this section is to show a criterion for an Ore extension of a regular multiplier Hopf coquasigroup to be a regular multiplier Hopf coquasigroup.
-First, let us introduce the definition of Ore extension of regular multiplier Hopf coquasigroups, which genelized the Ore extension of Hopf coquasigroups and multiplier Hopf algebras introduced in [3, 14].
+Let $(A,\Delta)$ denote a regular multiplier Hopf coquasigroup as defined in [14].
+The objective of this section is to establish a criterion for an Ore extension of a regular multiplier Hopf coquasigroup to also qualify as a regular multiplier Hopf coquasigroup.
+First, we shall introduce the definition of the Ore extension of regular multiplier Hopf coquasigroups, which extends the Ore extensions of Hopf coquasigroups and multiplier Hopf algebras as presented in [3, 16].
 Definition 3.1.
 Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups.
-The multiplier Hopf coquasigroup $(A[y;\tau,\delta],\Delta)$ is called the Ore extension of multiplier Hopf coquasigroups (or MHC-Ore extension in short) if $\Delta(y)=y\otimes r_{1}+r_{2}\otimes y$ for some $r_{1},r_{2}\in M(A)$.
-Remark 3.2. (1) As in [14], $y\otimes 1+r\otimes y$ in above definition belongs to $M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$.
-(2) If the coproduct $\Delta$ of multiplier Hopf coquasigroup $A$ is coassociative, then $A$ is a multiplier Hopf algebra and the Ore extension is actually the one introduced in [14].
-(3) If the underlying algebra of multiplier Hopf coquasigroup $A$ has an identity, then $A$ is a Hopf coquasigroup and the Ore extension is actually the one introduced in [3].
-By the definition of regular multiplier Hopf coquasigroup, for $y\in A[y;\tau,\delta]$, we have
-$(m\otimes\iota)(S\otimes\Delta)\Delta(y)=1_{M(A)}\otimes y=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(y),$
-$(\iota\otimes m)(\Delta\otimes S)\Delta(y)=y\otimes 1_{M(A)}=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(y),$
-that is
-$S(y)r_{1(1)}\otimes r_{1(2)}+S(r_{2})y\otimes r_{1}+S(r_{2})r_{2}\otimes y=1_{M(A)}\otimes y$
-$=yS(r_{1(1)})\otimes r_{1(2)}+r_{2}S(y)\otimes r_{1}+r_{2}S(r_{2})\otimes y,$
-$y\otimes S(r_{1})r_{1}+r_{2}\otimes S(y)r_{1}+r_{2(1)}\otimes S(r_{2(2)})y=y\otimes 1_{M(A)}$
-$=y\otimes r_{1}S(r_{1})+r_{2}\otimes yS(r_{1})+r_{2(1)}\otimes r_{2(2)}S(y).$
-Then
-$\hskip 40.00006pt{r_{1}}S({r_{1}})=S({r_{1}}){r_{1}}=1,\qquad{r_{2}}S({r_{2}})=S({r_{2}}){r_{2}}=1,$
-$S(y){r_{1(1)}}\otimes{r_{1(2)}}+S({r_{2}})y\otimes{r_{1}}=yS({r_{1(1)}})\otimes{r_{1(2)}}+{r_{2}}S(y)\otimes{r_{1}}=0,$
-${r_{2}}\otimes S(y){r_{1}}+{r_{2(1)}}\otimes S({r_{2(2)}})y={r_{2}}\otimes yS({r_{1}})+{r_{2(1)}}\otimes{r_{2(2)}}S(y)=0.$
-Let $r_{1}$ and $r_{2}$ be group-like elements in $M(A)$, then we have
-$\Delta(r_{i})=r_{i}\otimes r_{i},\quad\varepsilon(r_{i})=1,\quad S(r_{i})=r_{i}^{-1},\quad i=1,2,$
-$\big{(}S(y)r_{1}+S(r_{2})y\big{)}\otimes r_{1}=\big{(}yS(r_{1})+r_{2}S(y)\big{)}\otimes r_{1}=0,$
-$r_{2}\otimes\big{(}S(y)r_{1}+S(r_{2})y\big{)}=r_{2}\otimes\big{(}yS(r_{1})+r_{2}S(y)\big{)}=0.$
-By substituting the generator $y$ with $y^{\prime}=yr_{1}^{-1}$, it follows that the coproduct map satisfies $\Delta(y^{\prime})=y^{\prime}\otimes 1+r\otimes y^{\prime}$, where $r=r_{2}r_{1}^{-1}$ is a group-like element in the multiplier algebra $M(A)$.
-In the subsequent discussion, we consistently assume that the element $y$ in the Ore extension of a regular multiplier Hopf coquasigroup, the following relation holds:
-$\Delta\left(y\right)=y\otimes 1+r\otimes y$ (3.1)
-for some group-like element $r\in M(A)$. As usual, $A{d_{r}}\left(a\right)=raS\left(r\right)=ra{r^{-1}}$.
-Now we give the following properties for the Ore extension of a regular multiplier Hopf coquasigroup.
-Proposition 3.3. When $A[y;\tau,\delta]$ is a MHC-Ore extension of $A$, then
-$\varepsilon(y)=0,\quad S(y)=-r^{-1}y,\quad S^{-1}(y)=-\tau(r)^{-1}y,$
-$\hskip 60.00009pt\delta(r)=0,\quad\delta(r^{-1})=0.$
-proof. By applying the main property of counit, where $a=a_{(1)}\varepsilon(a_{(2)})$, it follows that $y=\varepsilon(y)+\varepsilon(r)y=\varepsilon(y)+y$, this implies $\varepsilon(y)=0$.
-By applying the main property of antipode, where $\varepsilon(a)=a_{(1)}S(a_{(2)})$, it follows that $0=\varepsilon(y)=y+rS(y)$, this implies $S(y)=-r^{-1}y$. Since
-$y=S^{-1}S(y)=S^{-1}(-r^{-1}y)=-S^{-1}(r^{-1}y)=-S^{-1}(y)S^{-1}(r^{-1})=-S^{-1}(y)r,$
-there is an element $a\in A$ such that $S^{-1}(y)=ay$. And applying the preceding formula we get
-$y=-ayr=-a\tau(r)y-a\delta(r),$
-this implies that $a=-\tau(r)^{-1}$ and $\delta(r)=0$, so we get that $S^{-1}(y)=-\tau(r)^{-1}y$. Since $\delta$ is a $\tau-derivation$, we have
-$0=\delta(1)=\delta(rr^{-1})=\delta(r)r^{-1}+\tau(r)\delta(r^{-1}),$
+The multiplier Hopf coquasigroup $(A[y;\tau,\delta],\Delta)$ is called an Ore extension of multiplier Hopf coquasigroup (or an MHC-Ore extension in short) if the following equation holds:
+$\Delta(y)=y\otimes 1_{M(A)}+r\otimes y$ (3.1)
+for a group-like element $r\in M(A)$.
+Remark 3.2.
+$(1)$ In what follows, we will simply write $1_{M(A)}$ as $1$ without causing any confusion.
+As stated in [16], the element $y\otimes 1+r\otimes y$ in equation $(\ref{3.1})$ belongs to $M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$.
+As usual, we denote $Ad_{r}(a)=raS(r)=rar^{-1}$.
+$(2)$ When the comultiplication $\Delta$ of the multiplier Hopf coquasigroup $A$ satisfies coassociativity, $A$ degenerates to a multiplier Hopf algebra and the Ore extension aligns with the one outlined in [16].
+As introduced in Section 2 of [16], all extensions depend on the algebraic structure of $A$, then extensions on multiplier Hopf coquasigroups are similar.
+$(3)$ If the underlying algebra of the multiplier Hopf coquasigroup AA possesses an identity element, $A$ can be classified as a Hopf coquasigroup, and the Ore extension corresponds to the one delineated in [3].
+Now, we present the properties associated with the Ore extension of a regular multiplier Hopf coquasigroup.
+Proposition 3.3. When $A[y;\tau,\delta]$ is an MHC-Ore extension of $A$, then
+$\varepsilon(y)=0,\quad S(y)=-r^{-1}y.$
+Proof. By applying the counit map $\varepsilon\otimes\iota$ to both sides of equation $(\ref{3.1})$, we obtain
+$y$ $=(\varepsilon\otimes\iota)\Delta(y)$
+$=(\varepsilon\otimes\iota)(y\otimes 1+r\otimes y)$
+$=\varepsilon(y)1+\varepsilon(r)y$
+$=\varepsilon(y)1+y.$
+Hence, it follows that $\varepsilon(y)=0$.
+Next, applying the same map $\varepsilon\otimes\iota$ again to equation $(\ref{3.1})$, we have
+$\varepsilon(y)$ $=m(\iota\otimes S)\Delta(y)$
+$=m(\iota\otimes S)(y\otimes 1+r\otimes y)$
+$=y+rS(y).$
 […]
-Next, we will present the main result of this section, which gives the necessary and sufficient conditions for the Ore extension of a multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup.
-Theoerm 3.4.
+We now proceed to state the principal theorem of this section, providing necessary and sufficient conditions under which the Ore extension of a multiplier Hopf coquasigroup again admits a multiplier Hopf coquasigroup structure.
+Theorem 3.4.
 Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups.
-Assume that $\tau$ is a surjective endomorphism on $A$.
-Then the regular multiplier Hopf coquasigroup $A[y;\tau,\delta]$ is a MHC-Ore extension of $A$ if and only if there is a group-like element $r\in M(A)$ such that the following conditions hold:
+Assume that $\tau$ is a surjective nondegenerate algebra endomorphism on $A$.
+Then the regular multiplier Hopf coquasigroup $A[y;\tau,\delta]$ is an MHC-Ore extension of $A$ if and only if there is a group-like element $r\in M(A)$ such that the following conditions hold:
 $(1)$ There exists a character $\chi$: $A\rightarrow k$ satisfying
-$\tau(a)=\chi(a_{(1)})a_{(2)};$ (3.2)
-$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)};$ (3.3)
+$\tau(a)=\chi(a_{(1)})a_{(2)}=(\chi\otimes\iota)\Delta(a)$ (3.2)
 for any $a\in A$.
-(2) The $\tau-derivation$ $\delta$ satisfies the relation
-$\Delta(\delta(a))=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})$ (3.4)
+$(2)$ The following equations hold
+$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}$ (3.3)
 for any $a\in A$.
-Proof. The proof will be conducted in three steps.
-step1. Comultiplication.
-Similar to the contruction in [14], assume that the comultiplication $\Delta$ can be extended from $A$ to $A[y;\tau.\delta]$ by $\Delta(y)=y\otimes 1+r\otimes y$.
+(3)(3)The $\tau$-derivation $\delta$ satisfies the relation
+$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes\iota)\Delta(a)+(r\otimes 1)\big{(}(\iota\otimes\delta)\Delta(a)\big{)}$ (3.4)
+for any $a\in A$.
+Proof. The proof proceeds in two parts, we begin with the necessity.
+1. Proof of necessity.
+Analogous to the construction in [16], the comultiplication $\Delta$ can be extended from $A$ to $A[y;\tau,\delta]$ by $\Delta(y)=y\otimes 1+r\otimes y$.
 Then the algebra homomorphism $\Delta$ preserves the relation $ya=\tau(a)y+\delta(a)$ for any $a\in A$, i.e.,
 $\Delta(y)\Delta(a)=\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}.$
-For any $a,b\in A$, we have that
+For any $b\in A$, we have
 $\Delta(y)\Delta(a)(1\otimes b)$
@@
 $=\tau(a_{(1)})y\otimes a_{(2)}b+\delta(a_{(1)})\otimes a_{(2)}b+ra_{(1)}\otimes\tau(a_{(2)}b)y+ra_{(1)}\otimes\delta(a_{(2)}b)$
-$=\tau({a_{(1)}})y\otimes{a_{(2)}}b+\delta({a_{(1)}})\otimes{a_{(2)}}b+r{a_{(1)}}\otimes\tau({a_{(2)}})\tau(b)y+r{a_{(1)}}\otimes\delta({a_{(2)}})b$
+$\overset{\bigstar}{=}\tau(a_{(1)})y\otimes a_{(2)}b+\delta(a_{(1)})\otimes a_{(2)}b+ra_{(1)}\otimes\tau(a_{(2)})\tau(b)y+r{a_{(1)}}\otimes\delta({a_{(2)}})b$
 $\quad+r{a_{(1)}}\otimes\tau({a_{(2)}})\delta(b)$
-$=\big{(}\tau({a_{(1)}})\otimes a_{(2)}\big{)}(1\otimes b)(y\otimes 1)+\big{(}r{a_{(1)}}{r^{-1}}\otimes\tau({a_{(2)}})\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)$
-$\quad+\big{(}r{a_{(1)}}{r^{-1}}\otimes\delta({a_{(2)}})\big{)}(r\otimes b)+\big{(}r{a_{(1)}}{r^{-1}}\otimes\tau({a_{(2)}})\big{)}\big{(}r\otimes\delta(b)\big{)}+\big{(}\delta(a_{(1)})\otimes a_{(2)}\big{)}(1\otimes b),$
-and
+$=\big{(}\tau(a_{(1)})\otimes a_{(2)}\big{)}(1\otimes b)(y\otimes 1)+\big{(}Ad_{r}(a_{(1)})\otimes\tau(a_{(2)})\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)$
+$\quad+\big{(}Ad_{r}(a_{(1)})\otimes\delta(a_{(2)})\big{)}(r\otimes b)+\big{(}Ad_{r}(a_{(1)})\otimes\tau(a_{(2)})\big{)}\big{(}r\otimes\delta(b)\big{)}$
+$\quad+\big{(}\delta(a_{(1)})\otimes a_{(2)}\big{)}(1\otimes b),$
+where the identity $ra_{(1)}\otimes\tau(a_{(2)}b)y=ra_{(1)}\otimes\tau(a_{(2)})\tau(b)y$ at step $\bigstar$ follows from the fact that $\tau$ is a nondegenerate algebra homomorphism. Furthermore,
 $\Big{(}\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}\Big{)}(1\otimes b)$
@@
 $=\Delta\big{(}\tau(a)\big{)}(y\otimes b)+\Delta\big{(}\tau(a)\big{)}(r\otimes yb)+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
-$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)y\big{)}+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
-$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}$
-$\quad+\Delta\big{(}\delta(a)\big{)}(1\otimes b).$
-So we have that $\Delta(y)\Delta(a)=\Delta(\tau(a))\Delta(y)+\Delta(\delta(a))$ holds if and only if the following formulas hold:
+$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)y\big{)}$
+$\quad+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
+$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)$
+$\quad+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b).$
+Therefore, the equality $\Delta(y)\Delta(a)=\Delta(\tau(a))\Delta(y)+\Delta(\delta(a))$ holds if and only if the following three conditions are satisfied: for all $b\in A$,
 $\Delta\big{(}\tau(a)\big{)}(1\otimes b)$ $=\tau(a_{(1)})\otimes a_{(2)}b,$
@@
 $\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$ $=ra_{(1)}\otimes\delta(a_{(2)}b)+\delta(a_{(1)})\otimes a_{(2)}b.$
-By the definition of $\delta\otimes\iota$ and $\iota\otimes\delta$ in [14, Lemma 2.5] and the extensions of $\tau\otimes\iota$, the above conditions change to the following:
-$\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)},$ (3.5)
+Using the properties of $\delta\otimes\iota$ and $\iota\otimes\delta$ from [16, Lemma 2.5], along with the extensions of $\tau\otimes\iota$, we reformulate the above conditions as follows:
+$\Delta\big{(}\tau(a)\big{)}=(\tau\otimes\iota)\Delta(a)=\tau(a_{(1)})\otimes a_{(2)},$ (3.5)
 $\Delta\big{(}\tau(a)\big{)}=A{d_{r}}(a_{(1)})\otimes\tau(a_{(2)}),$ (3.6)
-$\Delta\big{(}\delta(a)\big{)}=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)}).$
-[…]
-[…]
+$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes\iota)\Delta(a)+(r\otimes 1)\big{(}(\iota\otimes\delta)\Delta(a)\big{)}.$
+The final condition is specified by equation $(\ref{3.4})$.
+We now demonstrate that equations $(\ref{3.5})$ and $(\ref{3.6})$ imply the existence of a character $\chi$ such that both $(\ref{3.2})$ and $(\ref{3.3})$ are valid.
+Following the approach in the proof of Proposition 3.4 in [16], we define a map $\chi:A\rightarrow R(A)$ by
 $\chi(a)b=m(\iota\otimes S)\big{(}(b\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
-for any $a,b\in A$. Because $\tau$ is nondegenerate, for $b$ there are $a_{i}$ , $b_{i}$ such that $\sum{{a_{i}}}\tau\left({{b_{i}}}\right)=b$. Hence, the above definition is changed to the following:
-$\chi(a)b$ $=m(id\otimes S)\Big{(}\sum\big{(}a_{i}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}$
-$=m(id\otimes S)\Big{(}\sum(a_{i}\otimes 1)\big{(}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}$
-$=m(id\otimes S)\bigg{(}\sum(a_{i}\otimes 1)\Big{(}(\tau\otimes id)(b_{i}\otimes 1)\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}\bigg{)}$
-$=m(id\otimes S)\bigg{(}\sum(a_{i}\otimes 1)(\tau\otimes id)\Big{(}(b_{i}\otimes 1)\Delta\big{(}\tau(a)\big{)}\Big{)}\bigg{)}.$
-Hence $\chi(a)b\in A$ for any $a,b\in A$. In addition, for any $c\in A$,
-$\chi(a)(bc)$ $=m(id\otimes S)\big{(}(bc\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
-$=m(id\otimes S)\big{(}(b\otimes 1)(c\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
-$=b\big{(}m(id\otimes S)(c\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
+for any $a,b\in A$.
+Since $\tau$ is nondegenerate, for any $b$ there exist elements $a_{i}$ $b_{i}$ such that $\sum_{i=1}^{n}{{a_{i}}}\tau\left({{b_{i}}}\right)=b$.
+Hence, the expression above can be rewritten as:
+$\chi(a)b$ $=m(\iota\otimes S)\Big{(}\sum_{i=1}^{n}\big{(}a_{i}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes\iota)\Delta(a)\big{)}\Big{)}$
+$=m(\iota\otimes S)\Big{(}\sum_{i=1}^{n}(a_{i}\otimes 1)\big{(}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes\iota)\Delta(a)\big{)}\Big{)}$
+$=m(\iota\otimes S)\bigg{(}\sum_{i=1}^{n}(a_{i}\otimes 1)\Big{(}(\tau\otimes\iota)(b_{i}\otimes 1)\big{(}(\tau\otimes\iota)\Delta(a)\big{)}\Big{)}\bigg{)}$
+$=m(\iota\otimes S)\Big{(}\sum_{i=1}^{n}(a_{i}\otimes 1)(\tau\otimes\iota)\big{(}(b_{i}\otimes 1)\Delta(a)\big{)}\Big{)}.$
+It follows that $\chi(a)b\in A$ for all $a,b\in A$. In addition, for all $c\in A$,
+$\chi(a)(bc)$ $=m(\iota\otimes S)\big{(}(bc\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
+$=m(\iota\otimes S)\big{(}(b\otimes 1)(c\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
+$=b\big{(}m(\iota\otimes S)(c\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
 $=b\big{(}\chi(a)c\big{)}.$
-[…]
-$\chi(a)b=m(id\otimes S)\Big{(}(b\otimes 1)(\tau\otimes id)\Delta\big{(}\tau(a)\big{)}\Big{)}=b\varepsilon\big{(}\tau(a)\big{)},$
-and since $b$ is arbitrary, this proves that $\chi(a)=\varepsilon\big{(}\tau(a)\big{)}.$ Therefore, $\chi(A)\subseteq k$.
-We can regard $\chi$ as a mapping $\chi:A\rightarrow k$.
+This shows that $\chi(a)$ is a right multiplier on $A$ for any $a\in A$. According to the formula $(\ref{3.5})$, the properties of $S$ imply that for any $a,b\in A$
+$\chi(a)b\overset{(\ref{3.5})}{=}m(\iota\otimes S)\Big{(}(b\otimes 1)\Delta\big{(}\tau(a)\big{)}\Big{)}=b\varepsilon\big{(}\tau(a)\big{)},$
+this proves that $\chi(a)=\varepsilon\big{(}\tau(a)\big{)}$.
+Therefore, $\chi(A)\subseteq k$ and $\chi$ can be regarded as a map $\chi:A\rightarrow k$.
 Considering that $\tau$ is an algebra endomorphism and
@@
 $\chi(a+b)=\varepsilon\big{(}\tau(a+b)\big{)}=\varepsilon\big{(}\tau(a)\big{)}+\varepsilon\big{(}\tau(b)\big{)}=\chi(a)+\chi(b),$
-[…]
-$(id\otimes\Delta)\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)(1)}\otimes a_{(2)(2)}.$
-[…]
+which implies that $\chi$ is a character on $A$ and is nondegenerate.
+Now, for $\tau(a)\in A$, equation $(\ref{3.5})$ gives
+$(\iota\otimes\Delta)\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)(1)}\otimes a_{(2)(2)}.$
+Then by $(\ref{2.1})$ we have
 $S\big{(}\tau(a_{(1)})\big{)}a_{(2)(1)}\otimes a_{(2)(2)}=1\otimes\tau(a)=\tau(a_{(1)})S(a_{(2)(1)})\otimes a_{(2)(2)}.$ (3.7)
-By applying the formula $\chi\left(a\right)=\varepsilon\left(\tau\left(a\right)\right)$, $\tau$ can be derived from $\chi$,
+Using the formula $\chi(a)=\varepsilon\big{(}\tau(a)\big{)}$, we can express $\tau$ in terms of $\chi$,
 $\chi(a_{(1)})a_{(2)}$ $=\varepsilon\big{(}\tau(a_{(1)})\big{)}a_{(2)}=\varepsilon\big{(}\tau(a_{(1)})\big{)}\varepsilon(a_{(2)(1)})a_{(2)(2)}$
 $=\varepsilon\big{(}\tau(a_{(1)})\big{)}\varepsilon\big{(}S(a_{(2)(1)})\big{)}a_{(2)(2)}$
-$=\varepsilon\big{(}\tau(a_{(1)})\big{)}S(a_{(2)(1)})a_{(2)(2)}$
-$=\varepsilon(1)\tau(a)=\tau(a).$
-[…]
+$=\varepsilon\big{(}\tau(a_{(1)})S(a_{(2)(1)})\big{)}a_{(2)(2)}$
+$\overset{(\ref{3.7})}{=}\varepsilon(1)\tau(a)=\tau(a).$
+This establishes equation $(\ref{3.2})$. By substituting $\tau(a)=\chi(a_{(1)})a_{(2)}$(as established above) into both sides of equations $(\ref{3.5})$ and $(\ref{3.6})$, it ensues that
 $\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)},$
 $\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=A{d_{r}}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}.$
-[…]
-[…]
-$\Delta\big{(}\tau(a)\big{)}$ $=\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
-$=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}=\tau(a_{(1)})\otimes a_{(2)},$
-$\Delta\big{(}\tau(a)\big{)}$ $=\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
-$=Ad_{r}(a_{(1)})\otimes\chi(a_{(2)(1)})a_{(2)(2)}=Ad_{r}(a_{(1)})\otimes\tau(a_{(2)}).$
-[…]
-Therefore, we can obtain that the comultiplication of $A$ (i.e., the comultiplication on $M(A)$ ) can be extended to an algebra homomorphism $\Delta:M(A)[y;\tau,\delta]\rightarrow M(A)[y;\tau,\delta]\otimes M(A)[y;\tau,\delta]$, and $\Delta$ does not necessarily have to be coassociative.
+These two equations are exactly $(\ref{3.3})$.
+2. Proof of sufficiency.
+The sufficiency part of the proof is carried out in three steps.
+step 1. Comultiplication.
+Under conditions $(\ref{3.2})$-$(\ref{3.4})$, we proceed to show that $(\ref{3.5})$ and $(\ref{3.6})$ hold. We have
+$\Delta\big{(}\tau(a)\big{)}$ $\overset{(\ref{3.2})}{=}\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
+$\overset{(\ref{3.3})}{=}\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}\overset{(\ref{3.2})}{=}\tau(a_{(1)})\otimes a_{(2)},$
+$\Delta\big{(}\tau(a)\big{)}$ $\overset{(\ref{3.2})}{=}\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
+$\overset{(\ref{3.3})}{=}Ad_{r}(a_{(1)})\otimes\chi(a_{(2)(1)})a_{(2)(2)}\overset{(\ref{3.2})}{=}Ad_{r}(a_{(1)})\otimes\tau(a_{(2)}).$
+These equations confirm that $(\ref{3.5})$ and $(\ref{3.6})$ hold.
+Together with the necessity part of the proof, we conclude that $\Delta(y)\Delta(a)=\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}$ is satisfied.
+The comultiplication $\Delta:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$ is defined as above, note that it is not necessarily coassociative.
 Step 2. Counit.
-Assume that $A[y;\tau,\delta]$ already has the comultiplication defined as in the above step.
-In what follows, we will prove the existence of a counit denoted by $\overline{\varepsilon}$ extending $\varepsilon$ from $A$(in fact $M(A)$) to the Ore extension $M(A)[y;\tau,\delta]$.
-Define $\overline{\varepsilon}(y)=0$.
-We need to verify the equation
-$\overline{\varepsilon}(y)\varepsilon(x)=\varepsilon\big{(}\tau(x)\big{)}\overline{\varepsilon}(y)+\varepsilon\big{(}\delta(x)\big{)}$
-for any $x\in M(A)$, i.e., $\varepsilon\big{(}\delta(x)\big{)}=0$. In fact, we only have to prove the above formula is true for any $a\in A$ in the sense that
-$0=\varepsilon\big{(}\delta(xa)\big{)}=\varepsilon\big{(}\delta(x)\big{)}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\varepsilon\big{(}\delta(x)\big{)}.$
-[…]
-$\delta(a)$ $=m(id\otimes\varepsilon)\Delta\big{(}\delta(a)\big{)}$
-$=m(id\otimes\varepsilon)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
+Assume that $A[y;\tau,\delta]$ is equipped with the comultiplication defined in Step 1.
+Suppose further that the counit $\varepsilon$ can be extended to $A[y;\tau,\delta]$, with $\varepsilon(y)=0$, and remains an algebra homomorphism.
+Then the relation $ya=\tau(a)y+\delta(a)$ must be preserved under $\varepsilon$, i.e.,
+$\varepsilon(y)\varepsilon(a)=\varepsilon\big{(}\tau(a)\big{)}\varepsilon(y)+\varepsilon\big{(}\delta(a)\big{)}$
+for all $a\in A$. Substituting $\varepsilon(y)=0$, we find that such an extension exists if and only if
+$\varepsilon\big{(}\delta(a)\big{)}=0.$
+We now show that $\varepsilon\big{(}\delta(a)\big{)}=0$ indeed holds for all $a\in A$. Applying the map $m(\iota\otimes\varepsilon)$ to $(\ref{3.4})$, we obtain for any $a\in A$:
+$\delta(a)$ $=m(\iota\otimes\varepsilon)\Delta\big{(}\delta(a)\big{)}$
+$\overset{(\ref{3.4})}{=}m(\iota\otimes\varepsilon)\Big{(}(\delta\otimes\iota)\Delta(a)+(r\otimes 1)\big{(}(\iota\otimes\delta)\Delta(a)\big{)}\Big{)}$
+$=m(\iota\otimes\varepsilon)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
 $=\delta(a_{(1)})\varepsilon(a_{(2)})+ra_{(1)}\otimes\varepsilon\big{(}\delta(a_{(2)})\big{)}$
+$=\delta\big{(}a_{(1)}\varepsilon(a_{(2)})\big{)}+ra_{(1)}\otimes\varepsilon\big{(}\delta(a_{(2)})\big{)}$
 $=\delta(a)+ra_{(1)}\otimes\varepsilon\big{(}\delta(a_{(2)})\big{)}.$
-Therefore, $a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}=0$. Finally,
-$\varepsilon\big{(}\delta(a)\big{)}=\varepsilon\Big{(}\delta\big{(}\varepsilon(a_{(1)})a_{(2)}\big{)}\Big{)}=\varepsilon(a_{(1)})\varepsilon\big{(}\delta(a_{(2)})\big{)}=\varepsilon\Big{(}a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}\Big{)}=0$
-This proves $\varepsilon\big{(}\delta(a)\big{)}=0$ , for any $a\in A$. Furthermore, since
-$0=\varepsilon\big{(}\delta(xa)\big{)}$ $=\varepsilon\big{(}\delta(x)a+\tau(x)\delta(a)\big{)}$
-$=\varepsilon\big{(}\delta(a)\big{)}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\varepsilon\big{(}\delta(x)\big{)}$
-$=\varepsilon\big{(}\delta(x)\big{)}\Big{(}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\Big{)},$
-it follows that for any $x\in M(A)$, $\varepsilon\big{(}\delta(x)\big{)}=0$ holds true. Consequently, $\varepsilon$ can be extended to $M(A)[y;\tau,\delta]$.
-For all $\sum a_{i}y^{i},\sum b_{j}y^{j}\in A[y;\tau,\delta]$, it is easy to prove
-$(\overline{\varepsilon}\otimes\iota)\big{(}\Delta(\sum a_{i}y^{i})(1\otimes\sum b_{j}y^{j})\big{)}$ $=(\sum a_{i}y^{i})(\sum b_{j}y^{j}),$
-$(\iota\otimes\overline{\varepsilon})\big{(}(\sum b_{j}y^{j}\otimes 1)\Delta(\sum a_{i}y^{i})\big{)}$ $=(\sum b_{j}y^{j})(\sum a_{i}y^{i}).$
+Comparing both sides gives $a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}=0$. Finally,
+$\varepsilon\big{(}\delta(a)\big{)}=\varepsilon\Big{(}\delta\big{(}\varepsilon(a_{(1)})a_{(2)}\big{)}\Big{)}=\varepsilon(a_{(1)})\varepsilon\big{(}\delta(a_{(2)})\big{)}=\varepsilon\Big{(}a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}\Big{)}=0.$
+This establishes that $\varepsilon\big{(}\delta(a)\big{)}=0$, confirming the existence of an extension $\varepsilon:A[y;\tau,\delta]\rightarrow k$ that preserves the relation $ya=\tau(a)y+\delta(a)$.
+For any elements $\sum_{i=1}^{m}a_{i}y^{i},\sum_{j=1}^{n}b_{j}y^{j}\in A[y;\tau,\delta]$, it is straightforward to verify that
+$(\varepsilon\otimes\iota)\big{(}\Delta(\sum_{i=1}^{m}a_{i}y^{i})(1\otimes\sum_{j=1}^{n}b_{j}y^{j})\big{)}$ $=(\sum_{i=1}^{m}a_{i}y^{i})(\sum_{j=1}^{n}b_{j}y^{j}),$
+$(\iota\otimes\varepsilon)\big{(}(\sum_{j=1}^{n}b_{j}y^{j}\otimes 1)\Delta(\sum_{i=1}^{m}a_{i}y^{i})\big{)}$ $=(\sum_{j=1}^{n}b_{j}y^{j})(\sum_{i=1}^{m}a_{i}y^{i}).$
+Thus, we conclude that $\varepsilon:A[y;\tau,\delta]\rightarrow k$ indeed serves as a counit for $A[y;\tau,\delta]$.
 Step 3. Antipode.
-[…]
+Suppose $A[y;\tau,\delta]$ is endowed with the comultiplication and counit as defined above.
+If an antipode $S$ exists on $A[y;\tau,\delta]$, extended from the underlying algebra $A$ by setting $S(y)=-r^{-1}y$, then $S$ satisfies $(\ref{2.1})$ and $(\ref{2.2})$, while also preserving the relation $ya=\tau(a)y+\delta(a)$.
+This implies that
 $(m\otimes\iota)(S\otimes\Delta)\Delta(ya)=1\otimes ya=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(ya),$ (3.8)
@@
 and
-$S(a)S(y)=S(y)S\big{(}\alpha(a)\big{)}+S\big{(}\delta(a)\big{)}$ (3.10)
-[…]
-$1\otimes\delta(a)$ $=(m\otimes\iota)(S\otimes\Delta)\Delta\big{(}\delta(a)\big{)}$
+$S(a)S(y)=S(y)S\big{(}\tau(a)\big{)}+S\big{(}\delta(a)\big{)}.$ (3.10)
+We now proceed to prove equations $(\ref{3.8})$-$(\ref{3.10})$. For $\delta(a)\in A$ it follows from $(\ref{2.1})$ and $(\ref{3.4})$ that
+$1\otimes\delta(a)$ $\overset{(\ref{2.1})}{=}(m\otimes\iota)(S\otimes\Delta)\Delta\big{(}\delta(a)\big{)}$
+$\overset{(\ref{3.4})}{=}(m\otimes\iota)(S\otimes\Delta)\Big{(}(\delta\otimes\iota)\Delta(a)+(r\otimes 1)\big{(}(\iota\otimes\delta)\Delta(a)\big{)}\Big{)}$
 $=(m\otimes\iota)(S\otimes\Delta)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
@@
 $\quad+S(ra_{(1)})ra_{(2)(1)}\otimes\delta(a_{(2)(2)}).$ (3.11)
-Applying the formula mentioned above, we can obtain
+Applying the formula mentioned above, we deduce that for all $a,b\in A$
 $\big{(}S(b)\otimes 1\big{)}(m\otimes\iota)(S\otimes\Delta)\Delta(ya)$
@@
 $=(m\otimes\iota)(S\otimes\Delta)\bigg{(}\Big{(}\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}\Big{)}(b\otimes 1)\bigg{)}$
-$=(m\otimes\iota)(S\otimes\Delta)\Big{(}\tau({a_{(1)}})yb\otimes{a_{(2)}}+\tau({a_{(1)}})rb\otimes{a_{(2)}}y+\Delta\big{(}\delta(a)\big{)}(b\otimes 1)\Big{)}$
+$\overset{(\ref{3.5})}{=}(m\otimes\iota)(S\otimes\Delta)\Big{(}\tau(a_{(1)})yb\otimes a_{(2)}+\tau(a_{(1)})rb\otimes{a_{(2)}}y+\Delta\big{(}\delta(a)\big{)}(b\otimes 1)\Big{)}$
 $=(m\otimes\iota)\Big{(}S\big{(}\tau(a_{(1)})yb\big{)}\otimes a_{(2)(1)}\otimes a_{(2)(2)}+S\big{(}\tau(a_{(1)})rb\big{)}\otimes{a_{(2)(1)}}y\otimes a_{(2)(2)}$
@@
 $\quad+(m\otimes\iota)(S\otimes\Delta)\big{(}\delta(a_{(1)})b\otimes a_{(2)}+ra_{(1)}b\otimes\delta(a_{(2)})\big{)}$
-$=S\big{(}\tau({a_{(1)}})yb\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}+S\big{(}\tau({a_{(1)}})rb\big{)}{a_{(2)(1)}}y\otimes{a_{(2)(2)}}$
-$\quad+S\big{(}\tau({a_{(1)}})rb\big{)}{a_{(2)(1)}}r\otimes{a_{(2)(2)}}y+(m\otimes\iota)\Big{(}S\big{(}\delta(a_{(1)})b\big{)}\otimes{a_{(2)(1)}}\otimes{a_{(2)(2)}}$
-$\quad+S(r{a_{(1)}}b)\otimes\delta({a_{(2)(1)}})\otimes{a_{(2)(2)}}+S(r{a_{(1)}}b)\otimes r{a_{(2)(1)}}\otimes\delta({a_{(2)(2)}})\Big{)}$
+$\overset{(\ref{3.4})}{=}S\big{(}\tau(a_{(1)})yb\big{)}{a_{(2)(1)}}\otimes a_{(2)(2)}+S\big{(}\tau(a_{(1)})rb\big{)}a_{(2)(1)}y\otimes a_{(2)(2)}$
+$\quad+S\big{(}\tau({a_{(1)}})rb\big{)}a_{(2)(1)}r\otimes a_{(2)(2)}y$
+$\quad+(m\otimes\iota)\Big{(}S\big{(}\delta(a_{(1)})b\big{)}\otimes a_{(2)(1)}\otimes a_{(2)(2)}+S(ra_{(1)}b)\otimes\delta(a_{(2)(1)})\otimes{a_{(2)(2)}}$
+$\quad+S(ra_{(1)}b)\otimes ra_{(2)(1)}\otimes\delta(a_{(2)(2)})\Big{)}$
 $=S(b)S(y)S\big{(}\tau(a_{(1)})\big{)}a_{(2)(1)}\otimes a_{(2)(2)}+S(b)S(r)S\big{(}\tau(a_{(1)})\big{)}{a_{(2)(1)}}y\otimes a_{(2)(2)}$
-$\quad+S(b)S(r)S\big{(}\tau({a_{(1)}})\big{)}{a_{(2)(1)}}r\otimes{a_{(2)(2)}}y+S\big{(}\delta(a)b\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}$
+$\quad+S(b)S(r)S\big{(}\tau(a_{(1)})\big{)}a_{(2)(1)}r\otimes a_{(2)(2)}y+S\big{(}\delta(a_{(1)}b\big{)}a_{(2)(1)}\otimes a_{(2)(2)}$
 $\quad+S(ra_{(1)}b)\delta(a_{(2)(1)})\otimes a_{(2)(2)}+S(ra_{(1)}b)ra_{(2)(1)}\otimes\delta(a_{(2)(1)})$
-[…]
+$\overset{(\ref{3.7})}{=}S(b)S(y)\otimes\tau(a)+S(b)S(r)y\otimes\tau(a)+S(b)S(r)r\otimes\tau(a)y$
 $\quad+(S(b)\otimes 1)\Big{(}S\big{(}\delta(a_{(1)})\big{)}a_{(2)(1)}\otimes a_{(2)(2)}+S(ra_{(1)})\delta(a_{(2)(1)})\otimes{a_{(2)(2)}}$
 $\quad+S(ra_{(1)})ra_{(2)(1)}\otimes\delta(a_{(2)(2)})\Big{)}$
-=(3.11)${=}\big{(}S(b)S(y)+S(b)S(r)y\big{)}\otimes\tau(a)+S(b)\otimes\tau(a)y+(S(b)\otimes 1)(1\otimes\delta(a))$ )
+$\overset{(\ref{3.11})}{=}\big{(}S(b)S(y)+S(b)S(r)y\big{)}\otimes\tau(a)+S(b)\otimes\tau(a)y+(S(b)\otimes 1)(1\otimes\delta(a))$
 $=S(b)\otimes\tau(a)y+S(b)\otimes\delta(a)$
@@
 $=\big{(}S(b)\otimes 1\big{)}(1\otimes ya).$
-[…]
-[…]
+By comparing both sides of the equation above and using the nondegeneracy of the multiplication in $A$, we conclude that $(\ref{3.8})$ holds.
+Equation $(\ref{3.9})$ can be verified in a similar manner.
+To verify condition $(\ref{3.10})$, we make use of the identities $ya=\tau(a)y+\delta(a)$ and $S(y)=-r^{-1}y$ (where $r^{-1}y\in M(A[y;\tau,\delta])$), which lead to
 $-S(a)r^{-1}y$ $=-r^{-1}yS\big{(}\tau(a)\big{)}+S\big{(}\delta(a)\big{)}$
-$=-r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}y-r^{-1}\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}+S\big{(}\delta(a)\big{)},$
-for any $a\in M(A).$
-[…]
+$=-r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}y-r^{-1}\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}+S\big{(}\delta(a)\big{)}$
+for all $a\in A$. Rearranging terms, we obtain
+$\bigg{(}r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}-S(a)r^{-1}\bigg{)}y=-r^{-1}\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}+S\big{(}\delta(a)\big{)}.$
+It follows that condition $(\ref{3.10})$ is equivalent to the following pair of identities holding for all $a\in A$:
 $S(a)r^{-1}=r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)},$ (3.12)
 $\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}=rS\big{(}\delta(a)\big{)}.$ (3.13)
-for any $a\in M(A).$
-[…]
-In fact,
-$\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}$ $=\tau\Big{(}S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}\Big{)}=\tau\big{(}\chi({a_{(1)}})S({a_{(2)}})\big{)}=\chi({a_{(1)}})\tau\big{(}S({a_{(2)}})\big{)}$
-$=\chi({a_{(1)}})A{d_{r}}\Big{(}S(a_{(2)(2)})\chi\big{(}S(a_{(2)(1)})\big{)}\Big{)}$
-$=\chi(a_{(1)})S(a_{(2)(1)})Ad_{r}\big{(}S(a_{(2)(2)})\big{)}$
-$=Ad_{r}\big{(}S(a)\big{)}$
-[…]
-$S\big{(}\tau(a)\big{)}=S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}=\chi(a_{(1)})S(a_{(2)}).$
-[…]
-$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=rS\big{(}\delta(a)\big{)}.$ (3.14)
-[…]
-Applying $m(\iota\otimes S)$ to the equation
+It remains to show that $(\ref{3.12})$ and $(\ref{3.13})$ hold for every $a\in A$. Indeed,
+$\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}$ $\overset{(\ref{3.2})}{=}\tau\Big{(}S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}\Big{)}=\tau\big{(}\chi({a_{(1)}})S({a_{(2)}})\big{)}=\chi({a_{(1)}})\tau\big{(}S({a_{(2)}})\big{)}$
+$\overset{(\ref{3.2})(\ref{3.3})}{=}\chi({a_{(1)}})A{d_{r}}\Big{(}S(a_{(2)(2)})\chi\big{(}S(a_{(2)(1)})\big{)}\Big{)}$
+$=\chi({a_{(1)}})\chi\big{(}S(a_{(2)(1)})\big{)}A{d_{r}}\big{(}S(a_{(2)(2)})\big{)}$
+$=\chi\big{(}a_{(1)}S(a_{(2)(1)})\big{)}Ad_{r}\big{(}S(a_{(2)(2)})\big{)}$
+$\overset{(\ref{2.1})}{=}Ad_{r}\big{(}S(a)\big{)},$
+which confirms $(\ref{3.12})$.
+Proceeding to equation $(\ref{3.13})$, we utilize equation $(\ref{3.2})$, from which it follows that
+$S\big{(}\tau(a)\big{)}\overset{(\ref{3.2})}{=}S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}=\chi(a_{(1)})S(a_{(2)}).$
+Thus, equation $(\ref{3.13})$ can be reformulated equivalently as
+$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=rS\big{(}\delta(a)\big{)},$ (3.14)
+it suffices to demonstrate the validity of $(\ref{3.14})$. Applying the map $m(\iota\otimes S)$ to both sides of the identity
 $(b\otimes 1)\Delta\big{(}\delta(a)\big{)}=b\delta(a_{(1)})\otimes a_{(2)}+bra_{(1)}\otimes\delta(a_{(2)}),$
-it follows that
+we derive
 $m(\iota\otimes S)\Big{(}(b\otimes 1)\Delta\big{(}\delta(a)\big{)}\Big{)}=b\delta(a_{(1)})S(a_{(2)})+bra_{(1)}S(\delta\big{(}{a_{(2)}})\big{)}.$
-By applying the properties of the antipode, we have
+Using the defining property of the antipode, the left-hand side simplifies as
 $m(\iota\otimes S)\Big{(}(b\otimes 1)\Delta\big{(}\delta(a)\big{)}\Big{)}=b\varepsilon\big{(}\delta(a)\big{)}=0,$
-i.e,
+which implies
 $bra_{(1)}S\big{(}\delta(a_{(2)})\big{)}=-b\delta(a_{(1)})S(a_{(2)})\qquad\text{for any $a,b\in A$}.$
-And considering the nondegeneracy of the product of $A$, this means that the following equation holds for any $a\in A$
+By the nondegeneracy of the multiplication in $A$, we conclude that the following identity holds for all $a\in A$
 $a_{(1)}S\big{(}\delta(a_{(2)})\big{)}=-r^{-1}\delta(a_{(1)})S(a_{(2)}).$ (3.15)
-Hence, we have
-$rS\big{(}\delta(a)\big{)}$ $=rS(a_{(1)})a_{(2)(1)}S\big{(}\delta(a_{(2)(2)})\big{)}$
-=(3.15)${=}-rS(a_{(1)})r^{-1}\delta(a_{(2)(1)})S(a_{(2)(2)})$ )
-$=-Ad_{r}S({{a_{(1)}}})\delta(a_{(2)(1)})S(a_{(2)(2)}),$ (3.16)
-for any $a\in A$.
-In addition, from the action of $\delta$ on the element $\varepsilon(a)=a_{1}S(a_{(2)})$, it follows that
+Therefore, we have
+$rS\big{(}\delta(a)\big{)}$ $\overset{(\ref{2.1})}{=}rS(a_{(1)})a_{(2)(1)}S\big{(}\delta(a_{(2)(2)})\big{)}$
+$\overset{(\ref{3.15})}{=}-rS(a_{(1)})r^{-1}\delta(a_{(2)(1)})S(a_{(2)(2)})$
+$=-Ad_{r}\big{(}S(a_{(1)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$ (3.16)
+for all $a\in A$. Furthermore, applying $\delta$ to the identity $\varepsilon(a)=a_{(1)}S(a_{(2)})$, we find
 $0=\delta(a_{(1)})S(a_{(2)})+\tau(a_{(1)})\delta\big{(}S(a_{(2)})\big{)},$
-i.e.,
+which implies
 $\tau(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=-\delta(a_{(1)})S(a_{(2)}).$ (3.17)
-[…]
-$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}$ $=\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})a_{(2)(2)(1)}\big{)}\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
+By combining $(\ref{3.16})$ and $(\ref{3.17})$, we obtain the following
+$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}$ $\overset{(\ref{2.1})}{=}\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})a_{(2)(2)(1)}\big{)}\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
 $=\chi(a_{(1)})\tau\big{(}{S(a_{(2)(1)}})\big{)}\tau(a_{(2)(2)(1)})\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
-=(3.17)${=}-\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})\big{)}\delta(a_{(2)(2)(1)})S(a_{(2)(2)(2)})$ )
+$\overset{(\ref{3.17})}{=}-\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})\big{)}\delta(a_{(2)(2)(1)})S(a_{(2)(2)(2)})$
 $=-\tau\Big{(}S\big{(}\chi(a_{(1)})a_{(2)(1)}\big{)}\Big{)}\delta(a_{(2)(2)(1)})S(a_{(2)(2)(2)})$
-=(3.3)${=}-\tau\Big{(}S\big{(}\chi(a_{(1)(1)})a_{(1)(2)}\big{)}\Big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$ )
+$\overset{(\ref{3.3})}{=}-\tau\Big{(}S\big{(}\chi(a_{(1)(1)})a_{(1)(2)}\big{)}\Big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
 $=-\chi(a_{(1)(1)})\tau\big{(}S(a_{(1)(2)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
 $=-\chi(a_{(1)(1)})Ad_{r}\big{(}S(a_{(1)(2)(2)})\big{)}\chi\big{(}S(a_{(1)(2)(1)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
-$=-\chi(a_{(1)(1)})S(a_{(1)(2)(1)})Ad_{r}\big{(}S(a_{(1)(2)(2)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
+$=-\chi\big{(}a_{(1)(1)}S(a_{(1)(2)(1)})\big{)}Ad_{r}\big{(}S(a_{(1)(2)(2)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
 $=-Ad_{r}\big{(}S(a_{(1)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
-=(3.16)${=}rS\big{(}\delta(a)\big{)}$ )
-for any $a,b\in A$.
+$\overset{(\ref{3.16})}{=}rS\big{(}\delta(a)\big{)}$
 […]
+Remark 3.5.
+The proof of Theorem 3.4 adopts a strategy analogous to that of Theorem 3.3 in [3] and the theorem in [16], albeit with notable differences in technical execution.
+If the underlying algebra of $A$ is unital, then $A$ becomes a Hopf coquasigroup, recovering the framework studied in [3].
+And if the comultiplication in $A$ is coassociative, then $A$ reduces to a multiplier Hopf algebra, corresponding to the setting treated in [16].
+Example 3.6.
+$(1)$ Let $(A,\Delta,\varepsilon_{A},S_{A})$ be a multiplier Hopf algebra and $(H,\Delta,\varepsilon_{H},S_{H})$ a Hopf coquasigroup.
+Then the tensor product $A\otimes H$ forms a multiplier Hopf coquasigroup under the following structure:
+$(a\otimes h)(a^{\prime}\otimes h^{\prime})=aa^{\prime}\otimes hh^{\prime},$
+$\Delta(a\otimes h)=(a_{(1)}\otimes h_{(1)})\otimes(a_{(2)}\otimes h_{(2)})\in M\big{(}(A\otimes H)\otimes(A\otimes H)\big{)},$
+$\varepsilon_{A\otimes H}(a\otimes h)=\varepsilon_{A}(a)\varepsilon_{H}(h),$
+$S_{A\otimes H}(a\otimes h)=S_{A}(a)\otimes S_{H}(h),$
+where $a,a^{\prime}\in A$ and $h,h^{\prime}\in H$.
+If $A[y;\tau,\delta]$ is an Ore extension of $A$, then $A[y;\tau,\delta]\otimes H$ is an Ore extension of the multiplier Hopf coquasigroup $A\otimes H$ with
+$\Delta_{A\otimes H}(y\otimes 1_{H})=(y\otimes 1_{H})\otimes(1_{M(A)}\otimes 1_{H})+(r\otimes 1_{H})\otimes(y\otimes 1_{H}).$
+Similarly, $A\otimes H[y;\tau,\delta]$ is an Ore extension of $A\otimes H$ with $\Delta_{A\otimes H}(1_{M(A)}\otimes y)=(1_{M(A)}\otimes y)\otimes(1_{M(A)}\otimes 1_{H})+(1_{M(A)}\otimes r)\otimes(1_{M(A)}\otimes y)$.
+$(2)$ Let $A$ be an algebra spanned by elements $\{e_{p}e_{\alpha}\mid p\in\mathbb{Z},\alpha\in\mathbb{G}\}$ over $\mathbb{C}$, where $\mathbb{G}$ is a (finite or infinite) quasigroup, and the generators $e_{p}e_{\alpha}$ satisfy the orthogonality relations $(e_{p}e_{\alpha})(e_{q}e_{\beta})=\delta^{K}_{p,q}\delta^{K}_{\alpha,\beta}e_{p}e_{\alpha}$, with $\delta^{K}$ denoting the Kronecker delta.
+Define a comultiplication $\Delta$ on $A$ by
+$\Delta(e_{p}e_{\alpha})=\sum\limits_{\beta\in\mathbb{G}}\sum\limits_{k\in\mathbb{Z}}e_{k}e_{\beta}\otimes e_{p-k}e_{\beta^{-1}\alpha}.$
+Note that these infinite sums are well-defined in the strict topology on the multiplier algebra, i.e., multiplication with elements of $A$ yields finite sums.
+One may verify that $(A,\Delta)$ is a regular multiplier Hopf coquasigroup.
+The counit is given by $\varepsilon(e_{p}e_{\alpha})=\delta^{K}_{p,0}\delta^{K}_{\alpha,e}$, and the antipode by $S(e_{p}e_{\alpha})=e_{-p}e_{\alpha^{-1}}$.
+Following the construction in Example 3.9 of [16], choose a nonzero scalar $\lambda\in\mathbb{C}\setminus\{0\}$.
+Define $r\in M(A)$ as $r=\sum\limits_{k\in\mathbb{Z}}\lambda^{k}e_{k}$, a character $\chi:A\rightarrow\mathbb{C}$ by $\chi(e_{p}e_{\alpha})=\delta^{K}_{p+2,0}\delta^{K}_{\alpha,e}$, an algebra endomorphism $\tau(e_{p}e_{\alpha})=e_{p+2}e_{\alpha}$ and a $\tau$-derivation $\delta:A\rightarrow A$ by $\delta(e_{p}e_{\alpha})=0$.
+It is straightforward to verify that these definitions satisfy conditions (3.2)-(3.4).
+Consequently, the Ore extension $A[y;\tau,\delta]$ is established, thereby endowing AA with the structure of a multiplier Hopf coquasigroup in the sense of Theorem 3.4.
+At the end of this section, we introduce the following theorem concerning the *-structure of MHC-Ore extensions.
+In parallel with Definition 2.4 presented in [10], we first define the notion of a multiplier Hopf *-coquasigroup.
+Definition 3.7.
+If $A$ is a *-algebra over $\mathds{C}$, we call $\Delta$ a comultiplication if it is also a *-homomorphism.
+A multiplier Hopf *-coquasigroup is a *-algebra with a comultiplication, making it into a multiplier Hopf coquasigroup.
+Parallel to the proposition in [10], we have the following proposition for multiplier Hopf *-coquasigroups.
+Proposition 3.8. If $A$ is a multiplier Hopf *-coquasigroup, then for all $a\in A$, we have
+$(1)$ $\varepsilon(a^{*})=\varepsilon(a)^{-}$ (where $\varepsilon(a)^{-}$ denotes the conjugate of $\varepsilon(a)$);
+$(2)$ $S(S(a)^{*})^{*}=a$.
+Proof. $(1)$ Define $\tilde{\varepsilon}(a)=\varepsilon(a^{*})^{-}$. Consider the action of $(\tilde{\varepsilon}\otimes\iota)$ on $(1\otimes a)\Delta(b)$
+$(\tilde{\varepsilon}\otimes\iota)\big{(}(1\otimes a)\Delta(b)\big{)}$ $=\Big{(}(\varepsilon\otimes\iota)\big{(}\Delta(b^{*})(1\otimes a^{*})\big{)}\Big{)}^{*}=(b^{*}a^{*})^{*}=ab$
+$=(\varepsilon\otimes\iota)\big{(}(1\otimes a)\Delta(b)\big{)}.$
+This implies $\tilde{\varepsilon}=\varepsilon$.
+Substituting the definition of $\varepsilon$, we conclude that $\varepsilon(a)=\varepsilon(a^{*})^{-}$, and hence $\varepsilon(a^{*})=\varepsilon(a)^{-}$.
+$(2)$ Suppose $a\otimes b=\sum_{i=1}^{n}\Delta(a_{i})(1\otimes b_{i})$. Applying $m(S\otimes\iota)$ to both sides yields
+$S(a)b$ $=m(S\otimes\iota)(a\otimes b)=\sum_{i=1}^{n}m(S\otimes\iota)\big{(}\Delta(a_{i})(1\otimes b_{i})\big{)}$
+$=\sum_{i=1}^{n}\big{(}m(S\otimes\iota)\Delta(a_{i})\big{)}b_{i}=\sum_{i=1}^{n}\varepsilon(a_{i})b_{i}.$
+Thus, $S(a)b=\sum_{i=1}^{n}\varepsilon(a_{i})b_{i}$.
+Taking adjoints and applying the flip map gives $b^{*}\otimes a^{*}=\sum_{i=1}^{n}(b_{i}\otimes 1)\Delta^{cop}(a_{i}^{*})$.
+Next, applying $(\iota\otimes S^{-1})$ and multiplying on the right by $1\otimes c^{*}$, it follows from Lemma 3.1 in [15] that
+$b^{*}S^{-1}(a^{*})=\sum_{i=1}^{n}b_{i}^{*}\varepsilon(a_{i}^{*})c^{*}.$
+[…]
+Theorem 3.9. Let $A$ be a multiplier Hopf *-coquasigroup. Then the MHC-Ore extension $A[y;\tau,\delta]$ is also a multiplier Hopf *-coquasigroup provided the following conditions are satisfied:
+$(1)$ $\tau$ is an involution;
+$(2)$ $\delta(\tau+\iota)=0$ and $\delta*=*\delta$;
+$(3)$ $r^{*}=r$.
+Proof. We begin by extending the *-operator to $M(A[y;\tau,\delta])$ such that $y^{*}=y$. For $y\in M(A)[y;\tau,\delta]$ and any $a\in A$, conditions (1) and (2) imply
+$(ya)^{*}$ $=\big{(}\tau(a)y+\delta(a)\big{)}^{*}=y\tau(a)^{*}+\delta(a)^{*}$
+$=\tau\big{(}\tau(a)^{*}\big{)}y+\delta\big{(}\tau(a)^{*}\big{)}+\delta(a)^{*}=a^{*}y+\delta\big{(}\tau(a)\big{)}^{*}+\delta(a)^{*}$
+$=a^{*}y+\delta\big{(}\tau(a)+a\big{)}^{*}=a^{*}y+\big{(}\delta(\tau+\iota)(a)\big{)}^{*}$
+$=a^{*}y=a^{*}y^{*}.$
+Applying the result $a^{*}y=y\tau(a)^{*}+\delta(a)^{*}$ derived above, we have for any $a\in A$ and $x\in M(A)$
+$a^{*}x^{*}y$ $=(xa)^{*}y=y\tau(xa)^{*}+\delta(xa)^{*}$
+$=y(\tau(x)\tau(a))^{*}+(\delta(x)a+\tau(x)\delta(a))^{*}$
+$=y\tau(a)^{*}\tau(x)^{*}+a^{*}\delta(x)^{*}+\delta(a)^{*}\tau(x)^{*}$
+$=(y\tau(a)^{*}+\delta(a)^{*})\tau(x)^{*}+a^{*}\delta(x)^{*}$
+$=(a^{*}y)\tau(x)^{*}+a^{*}\delta(x)^{*}$
+$=a^{*}(y\tau(x)^{*}+\delta(x)^{*}),$
+this implies $x^{*}y^{*}=y^{*}\tau(x)^{*}+\delta(x)^{*}$.
+Hence, the involution * can be extended to a map $M(A)\to M(A)$ such that $y^{*}=y$.
+Moreover, for any element $\sum_{i=1}^{n}a_{i}y^{i}\in A[y;\tau,\delta]$, its image under the involution satisfies $(\sum_{i=1}^{n}a_{i}y^{i})^{*}\in A[y;\tau,\delta]$.
+It is straightforward to verify that $A[y;\tau,\delta]$ forms a *-algebra under the above definition.
+Next, we prove that $\Delta$ is a *-homomorphism. From condition (3), it follows that
+$\Delta(y)^{*}$ $=(y\otimes 1+r\otimes y)^{*}=y^{*}\otimes 1+r^{*}\otimes y^{*}$
+$=y\otimes 1+r\otimes y=\Delta(y)$
+$=\Delta(y^{*}).$
+Since $\Delta$ is an algebra homomorphism, it implies $\Delta(\sum_{i=1}^{n}a_{i}y^{i})^{*}=\Delta\big{(}(\sum_{i=1}^{n}a_{i}y^{i})^{*}\big{)}$.
+Furthermore, it is clear that $\varepsilon(y^{*})=0=\overline{\varepsilon(y)}$, and more generally, $\varepsilon\big{(}(\sum_{i=1}^{n}a_{i}y^{i})^{*}\big{)}=\overline{\varepsilon(\sum_{i=1}^{n}a_{i}y^{i})}$, so $\varepsilon$ is a *-homomorphism.
+It remains to prove that $S\big{(}S(y)^{*}\big{)}^{*}=y$. Using Proposition 3.3 and condition (3), we obtain
+$S(y)^{*}$ $=(-r^{-1}y)^{*}=-y^{*}(r^{-1})^{*}$
+$=-yr^{-1}=S(y).$
+and therefore
+$S\big{(}S(y)^{*}\big{)}$ $=S\big{(}S(y)\big{)}=S(-r^{-1}y)$
+$=-S(y)S(r^{-1})=r^{-1}yr$
+$=y.$
 […]
 4 Isomorphism of Ore extensions
-From Theorem 3.4, it follows that a MHC-Ore extension is determined by $\chi$, $r$, and $\delta$.
-Thus, $A(\chi,r,\delta)$ is used to denote the MHC-Ore extension of $A[y;\tau,\delta]$, where $\chi$ is a character of $A$, $r$ is a group-like element in $M(A)$ and $\delta$ is a $\langle\chi,r\rangle$-derivation.
-The corresponding Ore extension on the multiplier coquasigroup is denoted by $M(A)(\chi,r,\delta)$.
+According to Theorem 3.4, an Ore extension of a regular multiplier Hopf coquasigroup is uniquely determined by the parameters $\chi$, $r$, and $\delta$.
+We therefore denote such an Ore extension as $A[y;\tau,\delta]$ by $A(\chi,r,\delta)$, where $\chi$ is a character on $A$, $r$ is a group-like element in $M(A)$, and $\delta$ is a $\langle\chi,r\rangle$-derivation.
+The corresponding Ore extension at the level of multiplier coquasigroups is denoted by $M(A)(\chi,r,\delta)$.
 Now we can define an isomorphism of MHC-Ore extensions.
@@
 Two MHC-Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ of regular multiplier Hopf coquasigroups $A$ and $A^{\prime}$ are said to be isomorphic if there is an isomorphism of multiplier Hopf coquasigroups $\phi$: $A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ such that $\phi(A)=A^{\prime}$.
-Similarly to [14, Theorem 4.9], we have the following theorem for regular multiplier Hopf coquasigroups.
+Moreover, in the multiplier Hopf *-coquasigroup case, $A$ and $A^{\prime}$ are said to be isomorphic if furthermore $\phi$ is a *-homomorphism.
+In parallel with [16, Theorem 4.9], we have the following theorem for regular multiplier Hopf coquasigroups.
 Theorem 4.2.
-For multiplier Hopf coquasigroups $A$ and $A^{\prime}$, let their Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ be defined with coproducts: $\Delta(y)=y\otimes 1+r\otimes y$, $\Delta(y^{\prime})=y^{\prime}\otimes 1+r^{\prime}\otimes y^{\prime}$, where $r$ and $r^{\prime}$ are group-like elements in $A$.
+For multiplier Hopf coquasigroups $A$ and $A^{\prime}$, let their Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ be defined with communications: $\Delta(y)=y\otimes 1+r\otimes y$, $\Delta(y^{\prime})=y^{\prime}\otimes 1+r^{\prime}\otimes y^{\prime}$, where $r$ and $r^{\prime}$ are group-like elements in $A$.
 Then $A(\chi,r,\delta)$ is isomorphic to $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ if there is an isomorphism $\phi:A\rightarrow A^{\prime}$ such that
-$\phi(r)=r^{\prime},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},$ (4.1)
+$\phi(r)=r^{\prime},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},\quad\delta^{\prime}(\phi(a))=\phi(\delta(a))+\phi(\tau(a))d-d\phi(a),$
 where $d^{\prime}\in A^{\prime}$ such that $\Delta(d^{\prime})=d^{\prime}\otimes 1+r^{\prime}\otimes d^{\prime}$.
@@
 Let $\phi(y)=y^{\prime}+d^{\prime}$.
-Then $\phi$ can be extended from $\phi:A\rightarrow A^{\prime}$ to $\phi:A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$.
+Then $\phi$ extends from a map $\phi:A\rightarrow A^{\prime}$ to a homomorphism $\phi:A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$.
 We have
@@
 $=\tau^{\prime}\big{(}\phi(a)\big{)}y^{\prime}+\phi\big{(}\delta(a)\big{)}+\phi\big{(}\tau(a)\big{)}d^{\prime}-d^{\prime}\phi(a)+d^{\prime}\phi(a)$
-$=\phi\big{(}\tau(a)\big{)}y^{\prime}+\phi\big{(}\tau(a)\big{)}d^{\prime}+\phi\big{(}\delta(a)\big{)}=\phi\big{(}\tau(a)\big{)}(y^{\prime}+d^{\prime})+\phi\big{(}\delta(a)\big{)},$
-for any $a\in A$. Thus we get
+$=\phi\big{(}\tau(a)\big{)}y^{\prime}+\phi\big{(}\tau(a)\big{)}d^{\prime}+\phi\big{(}\delta(a)\big{)}$
+$=\phi\big{(}\tau(a)\big{)}(y^{\prime}+d^{\prime})+\phi\big{(}\delta(a)\big{)},$
+for any $a\in A$. Thus, we get
 $\phi(ya)=\phi(y)\phi(a).$
-Since
+Furthermore, we compute:
 $\Delta\big{(}\phi(y)\big{)}$ $=\Delta(y^{\prime}+d^{\prime})=\Delta(y^{\prime})+\Delta(d^{\prime})$
@@
 $=(y^{\prime}+d^{\prime})\otimes 1+r^{\prime}\otimes(y^{\prime}+d^{\prime}),$
-we can obtain
+which implies
 $\Delta\big{(}\phi(y)\big{)}=(\phi\otimes\phi)\Delta(y).$
-Considering the properties of $\varepsilon$, we obtain
-$d^{\prime}=(\varepsilon\otimes id)\big{(}\Delta(d^{\prime})\big{)}=\varepsilon(d^{\prime})+\varepsilon(r^{\prime})d^{\prime}=\varepsilon(d^{\prime})+d^{\prime},$
-this indicates $\varepsilon(d^{\prime})=0$. And by Proposition 3.3, we know that $\varepsilon(y)=0$ and $\varepsilon(y^{\prime})=0$. So we have
+Regarding the counit $\varepsilon$, we obtain
+$d^{\prime}=(\varepsilon\otimes\iota)\big{(}\Delta(d^{\prime})\big{)}=\varepsilon(d^{\prime})+\varepsilon(r^{\prime})d^{\prime}=\varepsilon(d^{\prime})+d^{\prime},$
+this indicates $\varepsilon(d^{\prime})=0$. By Proposition 3.3, we also know $\varepsilon(y)=0$ and $\varepsilon(y^{\prime})=0$. Hence,
 $\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y^{\prime}+d^{\prime})=\varepsilon(y^{\prime})+\varepsilon(d^{\prime})=0,$
-this means that $\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y)$.
-Considering the properties of $S$, we obtain
-$0=\varepsilon(d^{\prime})=m(S\otimes id)\Delta(d^{\prime})=S(d^{\prime})+S(r^{\prime})d^{\prime}=S(d^{\prime})+r{{}^{\prime-1}}d^{\prime},$
-this indicates $S(d^{\prime})=-r{{}^{\prime-1}}d^{\prime}$. Thus we have
-$\phi\big{(}S(y)\big{)}=\phi(-{r^{-1}}y)=-\phi({r^{-1}})\phi(y)=-r{{}^{\prime-1}}(y^{\prime}+d^{\prime}),$
-$S\big{(}\phi(y)\big{)}=S(y^{\prime}+d^{\prime})=S(y^{\prime})+S(d^{\prime})=-{r^{-1}}y^{\prime}-r{{}^{\prime-1}}d^{\prime}=-r{{}^{\prime-1}}(y^{\prime}+d^{\prime}),$
+which shows that $\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y)$.
+Now consider the antipode $S$, we find
+$0=\varepsilon(d^{\prime})=m(S\otimes\iota)\Delta(d^{\prime})=S(d^{\prime})+S(r^{\prime})d^{\prime}=S(d^{\prime})+r{{}^{\prime-1}}d^{\prime},$
+and therefore $S(d^{\prime})=-r{{}^{\prime-1}}d^{\prime}$. Consequently,
+[…]
+[…]
 it follows that $\phi\big{(}S(y)\big{)}=S\big{(}\phi(y)\big{)}$.
 […]
+Combining Theorem 3.9 and Theorem 4.2, we obtain the following corollary.
+Corollary 4.3.
+Let $A$ and $A^{\prime}$ be multiplier Hopf *-coquasigroups, with their *-structures denoted by *, the corresponding Hopf-Ore extensions for $A$ and $A^{\prime}$ are $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$.
+If there exists a multiplier Hopf *-algebra isomorphism $\phi:A\rightarrow A^{\prime}$ such that the hypotheses of Theorem 4.2 are satisfied (where $(d^{\prime})^{*}=d^{\prime}$), and the following conditions hold:
+(1) $\tau$ is an involution;
+(2) $\delta(\tau+\iota)=0$ and $\delta*=*\delta$;
+(3) $r^{*}=r$.
+Then $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ are isomorphic as Hopf-Ore extensions, and they are also isomorphic as multiplier Hopf *-coquasigroups.
+Proof. By Theorem 3.9 and Theorem 4.2, it suffices to show that $\phi$ preserves the *-operation, that is, for any $x\in A(\chi,r,\delta)$, we have $\phi(x^{*})=\phi(x)^{*}$.
+By the assumptions of this corollary, $\phi$ is a *-homomorphism on $A$, so for all $a\in A$, $\phi(a^{*})=\phi(a)^{*}$.
+Moreover, by Theorem 4.2, we have $\phi(y)=y^{\prime}+d^{\prime}$. It follows that
+$\phi(y^{*})=\phi(y)=y^{\prime}+d^{\prime},$
+$\phi(y)^{*}=(y^{\prime}+d^{\prime})^{*}=y^{\prime}+(d^{\prime})^{*}=y^{\prime}+d,$
+[…]
 Acknowledgements
-The work was partially supported by the China Postdoctoral Science Foundation (No. 2019M651764) and National Natural Science Foundation of China (No. 11601231).
+The work was partially supported by the Nanjing Agricultural University College Students’ Innovative Training Program (No.
+X2025103070154) China Postdoctoral Science Foundation (No.
+2019M651764) and National Natural Science Foundation of China (No.
+11601231).